### v1

# STG-Mamba: Spatial-Temporal Graph Learning via Selective State Space Model

Abstract
Spatial-Temporal Graph (STG) data is characterized as dynamic, heterogenous, and non-stationary, leading to the continuous challenge of spatial-temporal graph learning. In the past few years, various GNN-based methods have been proposed to solely focus on mimicking the relationships among node individuals of the STG network, ignoring the significance of modeling the intrinsic features that exist in STG system over time. In contrast, modern Selective State Space Models (SSSMs) present a new approach which treat STG Network as a system, and meticulously explore the STG system’s dynamic state evolution across temporal dimension. In this work, we introduce Spatial-Temporal Graph Mamba (STG-Mamba) as the first exploration of leveraging the powerful selective state space models for STG learning by treating STG Network as a system, and employing the Graph Selective State Space Block (GS3B) to precisely characterize the dynamic evolution of STG networks. STG-Mamba is formulated as an Encoder-Decoder architecture, which takes GS3B as the basic module, for efficient sequential data modeling. Furthermore, to strengthen GNN’s ability of modeling STG data under the setting of SSSMs, we propose Kalman Filtering Graph Neural Networks (KFGN) for adaptive graph structure upgrading. KFGN smoothly fits in the context of selective state space evolution, and at the same time keeps linear complexity. Extensive empirical studies are conducted on three benchmark STG forecasting datasets, demonstrating the performance superiority and computational efficiency of STG-Mamba. It not only surpasses existing state-of-the-art methods in terms of STG forecasting performance, but also effectively alleviate the computational bottleneck of large-scale graph networks in reducing the computational cost of FLOPs and test inference time.
1 Introduction
Spatial-temporal graph data is a kind of Non-Euclidean data which widely exists in our daily life, such as urban traffic network, metro system InFlow/OutFlow, social networks, regional energy load, weather observations, etc. Owning to the strong ability of capturing dynamic features across both temporal and spatial dimension, attention-based methods, such as Spatial-Temporal Transformer and its variants have recently been broadly implemented in spatial-temporal graph (STG) learning tasks. In the domain of STG traffic forecasting, seminal works have emerged that leverage these methods to enhance prediction accuracy. ASTGCN Guo et al. (2019) introduces an innovative approach combining spatial-attention & temporal attention mechanisms with graph neural networks. ADCT-Net Kong et al. (2024) presents to incorporate a dual-graphic cross-fusion Transformer for dynamic traffic forecasting. Huo et al. (2023) integrates hierarchical graph convolutional networks with Transformer networks to capture sophisticated STG relationships among traffic data. For weather forecasting, Liang et al. (2023) exemplifies how Transformer-based models can be utilized for accurate air quality forecasting across vast geographical areas. Chen et al. (2023) further demonstrates the power of GNNs in predicting air quality, emphasizing the importance of group-aware mechanisms in enhancing STG forecasting performance. In the realm of Social Recommendation, Huang et al. (2021) introduces a coupled GNN framework that incorporates inter-dependent knowledge graphs across both items and users. Wang et al. (2020a) explores the use of disentangled representations within a heterogeneous graph attention network to refine recommendation systems.
Although Transformer-based methods demonstrate notable improvements in STG relationship learning Chen et al. (2022); Cong et al. (2021), their modeling ability on large-scale STG networks and Long-Term Sequential Forecasting (LTSF) tasks is greatly hindered by the quadratic computational complexity $O(n^{2})$ provided by its core component–attention mechanism Zhou et al. (2021, 2022); Rampášek et al. (2022). Moreover, attention mechanisms usually approximate the full-length historical data, or encode all the context features, which may not be necessary for modeling the long-term dependencies of STG learning tasks. As such, the research world has been actively looking for counterparts of Attention-based models.
Lately, with the popularization of Mamba Gu and Dao (2023); Wang et al. (2024); Liu et al. (2024), Selective State Space Models (SSSM) have aroused considerable interest among researchers from computer vision and Natural Language Processing. SSSM is a variant of the State Space Models, which originated from Control Science and Engineering fields Lee et al. (1994); Friedland (2012) with strong theoretical foundation. State Space Models provide a professional framework to describe a physical system’s dynamics state evolution through sets of input, output, and state variables related by first-order differential or difference equations. The approach allows for a compact way to model and analyze systems with multiple inputs and outputs (MIMO) Aoki (2013).
As a pioneer deep learning-based SSSM, Mamba is proposed to address data-centric context learning for sequential tasks. Different from attention score calculation, Mamba construct selective state space models for the research objects, which encode contexts by using recurrent hidden states. The selection mechanism within Mamba allows it to control which part of the input data should be flowed into hidden states, as the significant context features that will influence the next-step embedding update.
STG learning can be treated as a complex process of understanding and forecasting the evolution of STG networks over time, which highly resembles the state space transition process where each state encapsulates the current configuration of a system and transitions represent changes over time. Deep learning-based SSSMs bring new horizons to STG learning tasks. However, great challenges pose in accurately and effectively adopting Mamba architectures for STG modeling, which is characterized as heterogeneous, dynamic, and non-stationary.
Motivated by the excellent long-term contextual modeling ability and low computational overhead of SSSMs, we propose Spatial-Temporal Graph Mamba (STG-Mamba) as the first deep learning-based SSSM for effective data-centric STG learning. The main contributions of this work are summarized in the following:
  • We make the first attempt to adapting SSSMs for STG learning tasks. A simple yet elegant way is employed to extend selective state space models to handle STG data. Specifically, we formulate the framework as Encoder-Decoder architecture, with $N$ stacked GS3B as basic modules and the proposed Kalman Filtering Graph Neural Network (KFGN) within it, to facilitate efficient Seq2Seq modeling.
  • A novel Graph Selective State Space Block (GS3B) is proposed for STG network pioneering integration with selective state space models, which performs input-dependent edge construction and dynamic node feature selection. GS3B is capable of capturing the sophisticated heterogenous dependencies among STG network, and accurately modeling the dynamic evolution of STG system over time, making it highly fitted for STG learning.
  • We introduce Kalman Filtering Graph Neural Network (KFGN) as the specialized adaptive spatiotemporal graph generation and upgrading method, which smoothly fits in within the SSSM-based context, ensuring the synchronous updates between graph structure and STG systems’s current state.
  • Extensive evaluations are carried on three open-sourced benchmark STG dataset. Results demonstrate that our STG-Mamba not only exceed other benchmark methods in terms of STG prediction performance, but also achieve $O(n)$ computational complexity, remarkably reducing the computational overhead compared with Transformers-based methods.
2 Preliminaries
Spatial-Temporal Graph (STG). STG is a kind of non-Euclidean structured data, which extensively existed in daily-life, business, and society. For a given graph network, if the node individuals have certain spatial connectivities at any given time step $\tau$, and the node features and/or graph structures are dynamically evolving over temporal axis, it could be defined as an STG network. Formally, we use $\mathbb{G}^{ST}=(V^{ST},E^{ST},A^{ST})$ to represent a given STG network. Here, $V^{ST}$ denotes the set of graph vertices (nodes) within the network, $E^{ST}$ denotes the edges of the graph network. $A^{ST}$ is the graph adjacency matrix, which indicates the calculated proximity between node individuals.
Spatial-Temporal Graph System (STG system). We for the first time define a network which consists of STG data as a Spatial-Temporal Graph System (STG system). In the theory foundation, a system is defined as the representation of a physical process, comprising state variables that describe the system’s current condition, input variables influencing the system’s state, and output variables reflecting the system’s response. For STG System, this framework is adapted to accommodate spatial dependencies and temporal evolution, structuring the system with nodes representing spatial entities, edges denoting spatial interactions, and state variables evolving over time to capture the dynamics of STG data.
State Space Models. State Space Models (SSMs) are a foundational concept in control science and system identification, offering a comprehensive framework to describe the dynamic evolutions of systems through mathematical theories. SSMs articulate a system’s behavior by defining a set of equations that represent the system’s state evolution over time, incorporating both control inputs and outputs.
Formally, SSMs are expressed through state equations $x_{t+1}=f(x_{t},u_{t})$ and observation equations $y_{t}=g(x_{t},u_{t})$, where $x_{t}$ denotes the state at time $t$, $u_{t}$ represents the control input, and $y_{t}$ is the observed output. The state transition function $f(\cdot)$ describes how the current state $x_{t}$ evolves to the next state $x_{t+1}$ under the influence of the control input $u_{t}$. $f(\cdot)$ captures the dynamics of the system, accounting for how internal and external factors modify the system’s state over time. The observation function $g(\cdot)$ is used for mapping the current state $x_{t}$ and control input $u_{t}$ to the observed output $y_{t}$. This function defines how the underlying state of the system is reflected in the measurements or observations collected.
As a variant of SSMs, Selective State Space Model emphasizes how to build the selection mechanism upon SSMs, which is highly resemble to the core idea of attention mechanism, making it a competitive counterpart of Transformer architecture.
STG Forecasting based on SSSMs. When employing Selective State Space Models (SSSMs) for spatial-temporal graph forecasting, the problem can be formulated as dynamically identifying and utilizing relevant portions of historical spatiotemporal data and graph structures to predict the future states of a STG system.
Given spatiotemporal graph $\mathbb{G}^{ST}=(V^{ST},E^{ST},A^{ST})$ and historical spatial-temporal sequential data $\rm{X}_{t-p+1:t}=\{X_{t-p+1},X_{t-p+2},...,X_{t}\}$, the objective is to utilize SSSMs to forecast future STG system’s states $\rm{\hat{X}_{t+1:t+k}}=\{\hat{X}_{t+1},...,\hat{X}_{t+k}\}$. This process is achieved by learning a mapping function $F_{SSSM}(\cdot)$ that dynamically selects relevant state transitions and interactions for prediction:
$F_{SSSM}(\rm{X}_{t-p+1:t};\mathbb{G}^{ST})=\rm{\hat{X}_{t+1:t+k}}$ (1)
3 Methodology
In this section, we start with introducing the model architecture of STG-Mamba and the formulation of STG Selective State Space Models. Following that, we elaborate the key contributions of this work, including the Kalman Filtering Graph Neural Networks (Section 3.2) which specifically designed for SSSMs to enable dynamic adaptive graph learning; and the Graph Selective State Space Blocks (GS3B) which acts as the basic modules for STG-Mamba (Section 3.3). Finally, we describe the computational efficiency of STG-Mamba in Section 3.4.
3.1 Formulation of STG Selective State Space Models
SSSMs like structured state space sequential models (S4) and Mamba, initially designed for sequential data, are adapted for Spatial-Temporal Graph (STG) data modeling by leveraging their structure to handle spatiotemporal dependencies dynamically. The adaptation involves discretizing the continuous system described by linear Ordinary Differential Equations (ODEs) into a format amenable to deep neural networks modeling and optimization, enhancing computational efficiency while capturing the evolving dynamics of the STG networks.
Figure 1: The comprehensive illustration of STG-Mamba’s Model Architecture.
Modern State Space Models all rely on a classical continuous system that maps a one-dimensional input function or sequence, denoted as $x(t)\in\mathbb{R}$, through intermediate implicit states $h(t)\in\mathbb{R}^{N}$ to an output $y(t)\in\mathbb{R}$. The aforementioned process can be represented as a linear Ordinary Differential Equation (ODE) as in Eq. 2:
$\begin{array}[]{lr}h^{\prime}(t)=\mathbb{A}h_{t}+\mathbb{B}x(t)\\ y(t)=\mathbb{C}h(t)\end{array}$ (2)
where $\mathbb{A}\in\mathbb{R}^{N\times N}$ represents the State Matrix, while $\mathbb{B}\in\mathbb{R}^{N\times 1},\mathbb{C}\in\mathbb{R}^{N\times 1}$ denote the Input Matrix and Output Matrix, respectively.
Obtaining the output sequence $y(t)$ at a given time-step $t$ necessitates determining $h(t)$ first, which poses great analytical challenges. As such, modern selective state space models address this challenge by converting the continuous system into a discretized version, enhancing the suitability for deep learning scenarios while boosting computational efficiency. A critical step in the discretization process involves the introduction of a time-scale factor $\Delta$, leading to the discrete representation of matrices $\mathbb{A}$ and $\mathbb{B}$ as $\overline{\mathbb{A}}$ and $\overline{\mathbb{B}}$, respectively, following a pre-determined discretization method, commonly the zero-order hold (ZOH) approach.
$\begin{array}[]{lr}\overline{\mathbb{A}}=\exp(\Delta\mathbb{A})\\ \overline{\mathbb{B}}=(\Delta\mathbb{A})^{-1}(\exp(\Delta A)-\mathbb{I})\cdot\Delta\mathbb{B}\end{array}$ (3)
After discretization, SSM-based methods can be computed in two ways: linear recurrent or global convolution, defined as in Eq. 4 and Eq. 5, respectively:
$\begin{array}[]{lr}h^{\prime}(t)=\overline{\mathbb{A}}h_{t}+\overline{\mathbb{B}}x(t)\\ y(t)=\mathbb{C}h(t)\end{array}$ (4)
$\begin{array}[]{lr}\overline{\mathbb{K}}=(\mathbb{C\overline{B}},\mathbb{C\overline{AB}},...,\mathbb{C\overline{A}^{L-1}\overline{B}})\\ y=x*\overline{\mathbb{K}}\end{array}$ (5)
where $\overline{K}\in\mathbb{R}^{L}$ denotes a structured convolutional kernel, and $L$ denotes the length of the input sequence $x$.
Figure 1 illustrates the comprehensive architecture of the proposed STG-Mamba. Specifically, we formulate the overall architecture following the Encoder-Decoder fashion for efficient sequential data modeling and prediction. In STG-Mamba, we leverage the Graph Selective State Space Block (GS3B) as the basic module, and repeat it $N$ times for Encoder and Decoder part. GS3B consists of several operations, including Layer Norm, Linear Layer, Kalman Filter Graph Neural Networks (KFGN), Graph State Space Layer (GSS-Layer), and Residual Connections. We specifically design KFGN and GSS-Layer for STG learning task build upon selective state space models.
3.2 Kalman Filtering Graph Neural Networks
Due to the dynamic, heterogenous, and non-stationary characteristics of STG data, the spatiotemporal dependencies of an STG system also evolve over time. When reflecting such graph dynamics in GNN formulation, it means the time-variant dynamic graph adjacency matrix. In existing literature, researchers usually employ various kinds of Attention-based methods to construct dynamic graph adjacency matrices, which may not be feasible in the context of state space model-based STG learning. On the other hand, Kalman Filtering (KF), as a well-known application of SSMs, has been extensively used in sequential data filtering and forecasting.
In light of that, we propose a novel approach which employs KF-based method to generate dynamic-evolving graphs, which delicately captures the system’s state-space transition for a given STG network. The method is termed as Kalman Filtering Graph Neural Networks (KFGN). To generate graph adjacency matrix $\mathbb{\alpha_{t}}$ at each time step using KFGN, the modeling procedure involves two steps as shown in the following:
Step 1: State Equation Calculation: Let $\mathbb{\alpha_{t}}$ denotes the graph adjacency matrix at time $t$, which indicates the current state of the STG system. The state equation is defined as $\mathbb{\alpha_{t}}=F_{t-1}\mathbb{\alpha_{t-1}}+W_{t-1}$, where $F_{t-1}$ is the state transition matrix that reflects changes in the graph adjacency matrix from $t-1$ to $t$, and $W_{t-1}$ represents the noise of the KF transition procedure.
Step 2: Observation Equation Calculation: The observed STG feature can be modeled as $X_{t}^{{}^{\prime}}=H_{t}\mathbb{\alpha_{t}}+V_{t}$, where $X_{t}^{{}^{\prime}}$ is the observation of STG feature at time $t$, $H_{t}$ is the observation matrix, and $V_{t}$ represents observation noise.
Iteratively applying these two equations through Kalman Filter enables dynamic estimation of the graph adjacency matrix $\alpha_{t}$ at each moment, adaptively capturing the evolving dynamics of the STG network. In practice, we employ a Neural Network-based method called Deep Kalman Filter (DKF) to automatically learn the state transition matrix $F_{t-1}$ and observation matrix $H_{t}$. Specifically, we adopt Gated Recurrent Unit (GRU) network as the architecture of Deep Kalman Filter. The GRU-based DKF is employed to learn the non-linear mapping for state transition and observation generation. The state transition matrix $F_{t-1}$ is modeled by GRU inner process to capture the non-linear relationships between states, predicting the state for the next moment. Similarly, the observation matrix $H_{t}$ is predicted under the current state by GRU. This process enables the system to automatically learn the dynamics of state transitions and observation generation from data without manually settings.
After receiving the $\alpha_{t}$, the next step of KFGN is the graph convolution operation, which is defined as:
$h_{t}^{kfgn}=(\omega_{gc}\odot\alpha_{t})x_{t}^{{}^{\prime}}$ (6)
where $\odot$ is the Hadamard product, and $x_{t}^{{}^{\prime}}$ denotes the input feature of KFGN at time $t$. $\omega_{gc}$ is the learnable weight matrix of KFGN.
3.3 GS3B: Graph Selective State Space Blocks
Inspired from Mamba, the GS3B is the major component of STG-Mamba, as depicted in Figure 1. After being processed by Layer Normalization, the input is divided into two sub-branches. In the first branch, the input STG embedding passed through a Linear Layer, followed by the proposed KFGN network and an activation function, before entering the GSS-Layer for selective feature learning. After that, the learned features are normalized by the Layer Norm. For the second branch, the input STG embedding undergoes a Linear Layer and an activation function. Subsequently, the two branches are fused together by an element-wise dot product operation. Finally, the merged embedding passes through a linear layer, and the outcome is combined with a residual connection to form the GS3B block’s output. In this research, Sigmoid Linear Unit (SiLU) is employed as the default activation function.
The GSS-Layer is a graph selective state space module specifically designed for STG data. GSS-Layer incorporates a selective mechanism which is capable of adjusting the SSM model’s parameters according to the input features. The characteristics enable the model to distinguish and retain key information which contributes significantly in STG learning tasks, while filtering out the irrelevant information. The pseudo-code of GSS-Layer (graph selective state space algorithm) is elaborated in Algorithm 1.
Algorithm 1 Pseudo Code of the Proposed Graph Selective State Space Algorithm
Input: $X\in\mathbb{R}^{batch\times L\times dim}$, the historical ground-truth STG data, where $batch$ is batch size of model training, $L$ is the historical time window length, $dim$ is the feature dimension of STG data;
$\mathbb{\alpha_{t}}$, the graph adjacency matrix of the current time step calculated by Deep Kalman Filter Network.
Params: $\mathbb{A}$ is the network param: nn.Parameter; $\mathbb{D}$ is the network param: nn.Parameter;
Operator: $Linear(\cdot)$, the linear projection layer;
Output:$y$, the processed feature with shape $\mathbb{R}^{batch\times L\times dim}$;
1:  $\Delta,\mathbb{B},\mathbb{C}=Linear(X),Linear(X),Linear(X)$
2:  $\overline{\mathbb{A}}=\exp(\Delta\mathbb{A})$
3:  $\overline{\mathbb{B}}=(\Delta\mathbb{A})^{-1}(\exp(\Delta\mathbb{A})-I)\cdot\Delta B$
4:  $h_{t}=(\overline{\mathbb{A}}\odot\mathbb{\alpha_{t}})h_{t-1}+\overline{\mathbb{B}}x_{t}$
5:  $y_{t}=\mathbb{C}h_{t}+\mathbb{D}x_{t}$
6:  $y=[y_{1},y_{2},...,y_{t},...,y_{L}]$.
7:  return $y$
3.4 Computational Efficiency of STG-Mamba
GS3B’s selection mechanism is introduced in Algorithm 1, which outlines the graph selective state space algorithm for efficiently processing STG data. It starts with initializing linear projections of the input STG data to obtain parameters $\Delta,\mathbb{B},\mathbb{C}$, which are used to compute the modified state transition matrix $\overline{\mathbb{A}}$ and input matrix $\overline{\mathbb{B}}$. These matrices incorporate the graph adjacency matrix $\mathbb{\alpha_{t}}$, updated by a Deep Kalman Filter Network, to calculate the hidden state $h_{t}$ for each time step. Finally, the output $y_{t}$ is generated by applying linear transformations to the hidden state and the input data, aggregating the results over the historical time window to form the output feature $y$. This algorithm effectively integrates dynamic graph structure information into SSSM, enhancing its capability to model sophisticated STG dynamics.
Leveraging a data-specific selection process, the increasing of sequence length $L$ or graph network node scale $N$ by $k$-fold can results in huge expansion of model parameters for Transformer-based methods. STG-Mamba introduces a sophisticated algorithm that is conscious of hardware capabilities, exploiting the layered structure of GPU memory to minimize the ensuing cost. In detail, for a batch size of $B$, STG feature dimension of $d^{\prime}$, and network node number of $N$, STG-Mamba reads the $O(BLd^{\prime}+Nd^{\prime})$ input data of $\mathbb{A},\mathbb{B},\mathbb{C},\Delta$ from High Bandwidth Memory (HBM), processes the intermediate stages sized $O(BLd^{\prime}N)$ in Static Random-Access Memory (SRAM), and dispatches the ultimate output, sized $O(BLd^{\prime})$, back to HBM. In contrast, if a Transformer-based model is employed, the HBM will read $O(BL^{2}d^{\prime})$ input data, and the computational size remains $O(BL^{2}d^{\prime})$ in SRAM. In common practices we often have $BL^{2}d^{\prime}\gg BLd^{\prime}+Nd^{\prime}$, which means SSSM-based approach can effectively decrease the computational overhead in large-scale parallel processing, such as neural network training. Additionally, by not retaining intermediate states, it also diminishes memory usage, since these states are recalculated during the gradient assessment in reverse phase. SSSM’s GPU-optimized design ensures STG-Mamba operates with a linear time complexity $O(L)$ relative to the length of the input sequence, marking a substantial speed advantage over traditional transformers’ dense attention computations, which exhibit quadratic time complexity $O(L^{2})$.
4 Experiments
4.1 Dataset Statistics and Baseline Methods
Datasets. For experimental evaluation, we meticulously select three real-world STG datasets from California road network speed records (PeMS04) Song et al. (2020), Hangzhou urban metro system entry/exit records (HZMetro) Liu et al. (2022), and weather station records across mainland China (KnowAir) Wang et al. (2020b). Acquired from the Caltrans Performance Measurement System (PeMS), PeMS04 records the average traffic speed from a total number of 307 sensors distributed on California Highway system. The time range of PeMS04 is 1st/Jan/2018 - 28th/Feb/2018, with 5-minute time interval. HZMetro was created with the In/Out transaction records of 80 stations of Hangzhou Metro system from 1st Jan 2019 to 25th Jan 2019, which takes 15-minute as data collection time interval. KnowAir documents weather observations in every 3-hour, encompassing a dataset that spans from 2015 to 2018. Within the KnowAir dataset, there are three subsets. In this study, we exclusively use subset No. 3, covering the period from 1/Sep/2016 to 31/Jan/2017. The detailed dataset statistics and descriptions are summarized in Table 1.
Table 1: Statistics and descriptions of the three datasets used for experimental evaluation.
Dataset PeMS04 HZMetro KnowAir
City California, USA Hangzhou, China 184 main cities across China
Data Type Network traffic speed Station InFlow/OutFlow Weather records
Total Nodes 307 80 184
Time Interval 5-minute 15-minute 3-hour
Time Range 1/Jan/2018 - 28/Feb/2018 1/Jan/2019 - 25/Jan/2019 1/Sep/2016 - 31/Jan/2017
Dataset Length 16,992 2,400 1,224
Baseline Methods. For fair comparison, we consider both benchmark spatiotemporal graph neural network (STGNN)-based methods, and Transformer-based methods for STG learning.
STGNN-based Methods:
(I) STGCN Yu et al. (2018): STGCN propose to capture the complex spatial-temporal dependencies within traffic network by graph neural networks, and design a fully convolutional approach for fast training.
(II) STSGCN Song et al. (2020): STSGCN presents a novel model which efficiently handling complex localized spatial-temporal correlations and data heterogeneity through a synchronous mechanism.
(III) STG-NCDE Choi et al. (2022): STG-NCDE leverages the concept of Neural Controlled Differential Equations (NCDEs) for spatial-temporal processing within a unified framework, achieving superior accuracy across six benchmark datasets.
(IV) DDGCRN Weng et al. (2023): DDGCRN dynamically captures the spatial-temporal features by separating the normal and abnormal signals and employing a novel graph convolutional recurrent network approach.
Attention (Transformer)-based Methods:
(V) ASTGCN Guo et al. (2019): ASTGCN leverages an attention-based spatial-temporal graph convolutional network to accurately predict STG data by dynamically capturing complex spatiotemporal correlations and patterns.
(VI) ASTGNN Guo et al. (2021): Attention based Spatial-Temporal Graph Neural Network incorporates a unique trend-aware self-attention and dynamic graph convolution to address temporal dynamics, spatial correlations, and the inherent periodicity and heterogeneity of STG data.
(VII) PDFormer Jiang et al. (2023): PDFormer introduces a tailored Transformer model, adept at capturing dynamic spatial dependencies and addressing the time delay in traffic condition propagation with its unique spatial self-attention and traffic delay-aware feature transformation.
(VIII) STAEformer Liu et al. (2023): STAEformer introduces a novel approach that enhances vanilla Transformer with a spatiotemporal adaptive embedding component, enabling it to capture complex STG patterns.
(IX) MultiSPANS Zou et al. (2024): MultiSPANS leverages multi-filter convolution modules, Transformers, and structural entropy optimization to effectively model complex multi-range dependencies in STG forecasting.
4.2 Implementation Settings
We split all the datasets with a ratio of 6:2:2 for Training/Validation/Testing sets along the time axis, respectively. Before starting model training session, all the data samples are normalized into range $[0,1]$ with MinMax normalization method. We set the batch size as 32 and employ AdamW as the Optimizer. The learning rate is initialized as $1e^{-3}$, with a weight decay of $5e^{-2}$. CosineAnnealingLR Loshchilov and Hutter (2016) is adopted as the learning rate scheduler, with a maximum of 50 iterations, and a minimum learning rate $1e^{-5}$. Model training epochs is set as 300. The number of Encoder/Decoder Layer is set as $N=8$ and $N^{\prime}=4$, respectively. We use the historical ground-truth data from the past 12 steps to forecast one future time step. Mean Squared Error (MSE) is employed as the Loss Function for optimization. And finally, we employ RMSE, MAE, MAPE as the evaluation metrics. All experiments are conducted under one NVIDIA A100 80GB GPU for neural computing acceleration.
4.3 Results Evaluation and Comparison
4.3.1 Experimental Results Comparison on Three STG Forecasting Tasks
A comprehensive evaluation between STG-Mamba and the proposed baseline methods is conducted, and we report the RMSE/MAE/MAPE performance results in Table 2. Our STG-Mamba consistently outperforms other benchmark methods on the three tasks, except for the MAPE criteria of PeMS04 (Flow). In PeMS04 (Flow) forecasting task, STAEformer shows slightly better MAPE results than STG-Mamba. It can also be easily observed that Transformer-based methods which integrate GNNs inside the architecture always gain superior performance than sole GNN-based methods. For instance, PDFormer, STAEformer, and MultiSPANS always show better performance than STGCN, STSGCN, and STG-NCDE. The reason is that although GNN is an effective approach of modeling STG tasks, Transformer-based methods have stronger ability to capture local and long-term dynamic dependencies. Within the GNN-based methods, STG-NCDE and DDGCRN stand out for their competitive performance. STG-NCDE models continuous-time dynamics within STG networks through neural controlled differential equations (NCDE), offering a more precise representation of spatiotemporal feature changes over time compared to other discrete GNN-based methods.
As a new contender poised to challenge Transformer architecture, Mamba introduces an innovative approach by replacing the foundational Attention mechanism with modern selective state space model (SSSM). SSSM diverges from Attention by utilizing a continuous-time model that captures the dynamic evolution of spatial-temporal dependencies more naturally and efficiently. Unlike Attention, which computes weights across all pairs of input and output positions and can be computationally intensive, Mamba’s selective state space model directly captures the temporal evolution of features, enabling it to scale more gracefully with the sequence length and complexity of spatial-temporal data. Furthermore, we facilitate STG-Mamba with Deep Kalman Filtering-based adaptive spatiotemporal graph evolution, and the elaborately designed Graph Selective State Space Block (GS3B), making it inherit the advantages of modern SSSM, and highly suitable for STG learning tasks.
Table 2: Performance Evaluation and Comparison with Benchmark Baselines on PeMS04, HZMetro, KnowAir Datasets.
Model PeMS04 (Flow) HZMetro KnowAir
RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE
STGCN 35.55 22.70 14.59 34.85 21.33 13.47 20.64 15.27 12.92
STSGCN 33.65 21.19 13.90 33.24 20.79 13.06 20.25 15.22 12.69
STG-NCDE 31.09 19.21 12.76 32.91 20.75 12.88 20.08 14.93 11.87
DDGCRN 30.51 18.45 12.19 31.69 19.52 12.45 19.63 12.45 11.84
ASTGCN 35.22 22.93 16.56 34.36 21.12 13.50 20.37 15.07 12.11
ASTGNN 31.16 19.26 12.65 32.75 20.63 12.47 19.96 12.38 11.23
PDFormer 29.97 18.32 12.10 30.18 19.13 11.92 18.65 12.21 11.06
STAEformer 30.18 18.22 11.98 29.94 18.85 12.03 18.49 11.93 10.75
MultiSPANS 30.46 19.07 13.29 30.31 18.97 11.85 18.57 12.24 10.88
STG-Mamba 29.53 18.09 12.12 29.26 18.38 11.67 18.05 11.73 10.66
4.3.2 Computational Efficiency Evaluation
STG-Mamba offers substantial improvement in computational efficiency in addition to performance gain. In order to quantitively evaluate computational efficiency, we select the widely recognized Floating Point Operations (FLOPs) and Inference Time as criteria. For both of the two criteria, a smaller value indicates better model performance. STG-Mamba is compare with other Transformer-based benchmark methods: ASTGNN, PDFormer, and STAEformer, on the same testing set.
Table 3 showcases the inference time comparison on PeMS04 and KnowAir testing set. It can be seen that ASTGNN is relatively redundancy, with the lowest inference time among the compared Transformer-based methods. STG-Mamba gains 18.09% and 20.54% performance increase in inference time than the 2nd best performance model (STAEformer).
We further compare the computational FLOPs increment with accordance to the scale complexity of STG network. Specifically, we use the number of node individuals to represent the scale of STG network. In the experiments, we select different number of nodes from PeMS04 and KnowAir. We select $[50,100,150,200,250,300]$ and $[30,60,90,120,150,180]$ node individuals for PeMS04 and KnowAir dataset, respectively. STG-Mamba is compared with the benchmark STAEformer in every settings. Figure 2 visualizes the FLOPs evaluation under different STG scale complexity. We can easily find that STG-Mamba enjoys a linear increment in FLOPs, while STAEformer shows quadratic increment in FLOPs. At the initial small-scale setting, the FLOPs for STG-Mamba/STAEformer are quite similar. However, with the increasing of network complexity and scale, the difference becomes more significant. At the 300 node setting of PeMS04, the FLOPs for STG-Mamba and STAEformer is 36.85G and 92.49G (2.50 times more), respectively. At the 180 node setting of KnowAir, the FLOPs for STG-Mamba and STAEformer is 57.08G and 118.13G (2.07 times more), respectively.
Benefiting from the linear computational increment brought by modern selective state space models, STG-Mamba and alike models demonstrate stronger capability and computational efficiency for modeling large-scale STG data, which holds promise for addressing the computational overhead in future large-scale STG learning.
Table 3: Test inference time comparison with benchmark Transformer-based approaches.
Model Inference Time on PeMS04 Inference Time on KnowAir
ASTGNN 4.85 s 5.31s
PDFormer 3.93 s 4.28s
STAEformer 3.76 s 4.09s
STG-Mamba 3.08 s 3.25s
performance $\uparrow$ +18.08% +20.54%
Figure 2: (Computational Efficiency Eval) FLOPs comparison between STG-Mamba and STAEformer on PeMS04/KnowAir dataset under different STG node number settings.
4.4 Ablation Study
Figure 3: Model component analysis of STG-Mamba.
To investigate the effectiveness of each model component within STG-Mamba, we further design three kinds of model variants, and evaluate their forecasting performance on PeMS04, HZMetro, and KnowAir dataset. The detailed description of these model variants are introduced in the following:
  1. 1.
    STG-Mamba: The full STG-Mamba model without any modification.
  2. 2.
    STG-Mamba w/o KFGN: We replace the proposed Kalman Filtering Graph Neural Networks (KFGN) within STG-Mamba with the standard Graph Convolution Network (GCN).
  3. 3.
    STG-Mamba w/o GSS-Layer: The GSS-Layer with proposed Graph Selective State Space Algorithm is replaced by the standard selective state space module as introduced in Gu and Dao (2023).
  4. 4.
    Transformer w KFGN: The GS3B block is reformulated as Transformer block (similar to ASTGNN Guo et al. (2021)), which consists of attention mechanism and the proposed KFGN.
As illustrated in Figure 3, STG-Mamba w/o KFGN shows slight worse performance than the full model, demonstrating the necessity and suitability of employing Kalman filtering-based dynamic GNN for SSSM-based methods. Different from the mainstream methods which employ various attention mechanisms to upgrade STG network’s adjacency matrix, our KFGN is more consistent with the SSSM context, and at the same time keep the linear algorithm complexity. STG-Mamba w/o GSS-Layer performs even worse than w/o KFGN. This is because in our design, GSS-Layer will take the generated STG adjacency matrix $\alpha_{t}$ at this timestamp as input for selective state space algorithm’s better upgradation. Replacing GSS-Layer with the standard selective state space layer leads to a less suitable version for STG learning. Finally, when we change from Mamba-based architecture to Transformer-based architecture, there is a substantial performance drop among all the three tasks. This discovery proves the effectiveness of modern SSSM over attention-based designs.
5 Conclusions
In this work, we first propose a novel spatial-temporal graph (STG) learning model named STG-Mamba that leverages modern selective state space models (SSSM) for accurate and efficient STG forecasting tasks. STG-Mamba is formulated as Encoder-Decoder structure, with the Graph Selective State Space Block (GS3B) as basic module. GS3B’s selection method facilitates input-dependent edge construction and node feature selection, successfully integrates STG network with SSSMs. Through the Kalman Filtering Graph Neural Networks (KFGN), the STG network achieves a smooth and dynamic evolution process of spatiotemporal graph adjacency matrix upgrading, which align with the whole graph selective state space modeling process. Compared with Attention-based methods, our STG-Mamba is able to achieve linear-time complexity and substantial reduction in FLOPs and test inference time. We conduct extensive empirical studies on three open-sourced STG forecasting datasets from urban traffic, metro In/Out flow, and weather records. Experimental results demonstrate the consistent superiority of STG-Mamba over other benchmark methods in both forecasting performance and computational efficiency. We believe that STG-Mamba presents a brand new promising approach to general STG learning fields, offering competitive model performance and input-dependent context-aware feature selection under an affordable computational cost.

### v2

@@
 3.3 GS3B: Graph Selective State Space Blocks
+Figure 2: The selection mechanism of STG-Mamba that facilitates input-dependent feature selection and focusing.
 Inspired from Mamba, the GS3B is the major component of STG-Mamba, as depicted in Figure 1.
@@
 In this research, Sigmoid Linear Unit (SiLU) is employed as the default activation function.
+We further comprehend the detailed process of the STG-Mamba’s Selection Mechanism in Figure 2.
 The GSS-Layer is a graph selective state space module specifically designed for STG data.
@@
 In contrast, if a Transformer-based model is employed, the HBM will read $O(BL^{2}d^{\prime})$ input data, and the computational size remains $O(BL^{2}d^{\prime})$ in SRAM.
-In common practices we often have $BL^{2}d^{\prime}\gg BLd^{\prime}+Nd^{\prime}$, which means SSSM-based approach can effectively decrease the computational overhead in large-scale parallel processing, such as neural network training.
+In common practices we often have $BL^{2}d\gg BLd^{\prime}+Nd^{\prime}$, which means SSSM-based approach can effectively decrease the computational overhead in large-scale parallel processing, such as neural network training.
 Additionally, by not retaining intermediate states, it also diminishes memory usage, since these states are recalculated during the gradient assessment in reverse phase.
@@
 Model training epochs is set as 300.
-The number of Encoder/Decoder Layer is set as $N=8$ and $N^{\prime}=4$, respectively.
+The number of Encoder/Decoder Layer is set as $N=4$ and $N^{\prime}=4$, respectively.
 We use the historical ground-truth data from the past 12 steps to forecast one future time step.
@@
 STG-NCDE models continuous-time dynamics within STG networks through neural controlled differential equations (NCDE), offering a more precise representation of spatiotemporal feature changes over time compared to other discrete GNN-based methods.
+Figure 3 presents a direct performance visualization comparison among STG-Mamba and other superior Transformer-based baselines (ASTGNN/PDFormer/STAEformer) on PeMS04, HZMetro, and KnowAir dataset, respectively.
+We visualize the predictions over 72 consecutive time steps using the average results of multiple 12-step forecast settings.
+For each dataset, four node individuals are randomly selected (i.e., 4 random traffic sensors of PeMS04, 4 random metro stations of HZMetro, and 4 random weather stations of KnowAir).
+The results indicate that STG-Mamba consistently yields the prediction curve closest to the ground truth waveform, featuring more accurate STG sequence peak values and fewer fluctuations.
+Figure 3: STG forecasting results visualization and comparison among STG-Mamba and other superior Transformer-based baselines.
 As a new contender poised to challenge Transformer architecture, Mamba introduces an innovative approach by replacing the foundational Attention mechanism with modern selective state space model (SSSM).
@@
 STG-Mamba offers substantial improvement in computational efficiency in addition to performance gain.
-In order to quantitively evaluate computational efficiency, we select the widely recognized Floating Point Operations (FLOPs) and Inference Time as criteria.
+In order to quantatively evaluate computational efficiency, we select the widely recognized Floating Point Operations (FLOPs) and Inference Time as criteria.
 For both of the two criteria, a smaller value indicates better model performance.
@@
 Figure 4: (Computational Efficiency Eval) FLOPs comparison between STG-Mamba and STAEformer on PeMS04/KnowAir dataset under different STG node number settings.
+4.4 Robustness Analysis
+Table 4: Robustness analysis on PeMS04. We compare the forecasting performance among STG-Mamba and 4 superior baseline methods during rush/non rush hours, weekend/non weekend settings.
+Models Rush Hours Non Rush Hours Weekend Non Weekend
+RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE
+ASTGNN 33.57 21.05 13.81 29.76 18.22 12.49 30.66 19.08 13.35 31.82 19.53 12.28
+PDFormer 31.69 19.54 12.38 28.95 18.16 11.97 29.87 18.26 12.04 30.39 18.68 12.23
+STAEformer 31.36 19.10 12.22 29.47 17.98 11.93 29.83 18.19 11.96 30.26 18.42 12.27
+MultiSPAN 30.95 19.28 13.31 30.04 18.37 12.54 30.13 18.42 12.73 30.52 19.31 13.34
+STG-Mamba 29.58 18.13 12.15 29.23 18.04 11.91 29.31 18.05 11.92 29.57 18.15 12.16
+STG data exhibit notable periodicity and diversity.
+For urban mobility/traffic data, there are distinct differences in traffic characteristics between the peak travel periods in the morning and evening and the none-peak periods.
+Moreover, the traffic patterns over the 24 hours of weekdays and those of weekends also differ substantially.
+In light of the variations in STG characteristics due to external environmental changes, it is imperative to assess whether deep learning models can accurately and effectively model the spatial-temporal dependencies under various external conditions.
+As such, we establish four distinct external scenarios: (a) Traffic rush hours: Monday to Friday, 8:00-11:00 AM and 16:00-19:00 PM; (b) Non-rush hours: encompassing the remaining time periods from Monday to Friday; (c) Weekend: including all hours on Saturday and Sunday; and (d) Non-weekend: including all time periods from Monday to Friday.
+Extensive experiments were conducted on the PeMS04 dataset.
+Table 4 presents the STG forecasting results corresponding to the four traffic scenarios.
+We compare the prediction performance of STG-Mamba with the other four superior baselines, and report the RMSE/MAE/MAPE metrics.
+An ideal deep learning-based STG system should maintain robustness to external disturbances (changes of external environment or events).
+This is specifically reflected in the fact that when facing different scenarios, the model performance should maintain as little difference as possible, rather than being good at some scenarios but being bad at other occasions.
+Among the compared methods, ASTGNN, PDFormer, and STAEformer exhibit relatively large differences in the prediction results between the Rush hour and Non-rush Hour scenarios.
+In the Weekend/Non Weekend scenario, although the performance differences of each model are not as significant as in the Rush hour/Non-Rush hour scenario, we observe that STG-Mamba still shows the smallest performance difference (the maximum robustness) and achieves the best RMSE/MAE/MAPE metrics.
+In summary, STG-Mamba proves the best robustness and outperforms existing baseline methods even in face of complex and diversity traffic scenes.
 4.5 Ablation Study

### v3

@@
 In contrast, modern Selective State Space Models (SSSMs) present a new approach which treat STG Network as a system, and meticulously explore the STG system’s dynamic state evolution across temporal dimension.
-In this work, we introduce Spatial-Temporal Graph Mamba (STG-Mamba) as the first exploration of leveraging the powerful selective state space models for STG learning by treating STG Network as a system, and employing the Graph Selective State Space Block (GS3B) to precisely characterize the dynamic evolution of STG networks.
-STG-Mamba is formulated as an Encoder-Decoder architecture, which takes GS3B as the basic module, for efficient sequential data modeling.
-Furthermore, to strengthen GNN’s ability of modeling STG data under the setting of SSSMs, we propose Kalman Filtering Graph Neural Networks (KFGN) for adaptive graph structure upgrading.
-KFGN smoothly fits in the context of selective state space evolution, and at the same time keeps linear complexity.
+In this work, we introduce Spatial-Temporal Graph Mamba (STG-Mamba) as the first exploration of leveraging the powerful selective state space models for STG learning by treating STG Network as a system, and employing the Spatial-Temporal Selective State Space Module (ST-S3M) to precisely focus on the selected STG latent features.
+Furthermore, to strengthen GNN’s ability of modeling STG data under the setting of selective state space models, we propose Kalman Filtering Graph Neural Networks (KFGN) for dynamically integrate and upgrade the STG embeddings from different temporal granularities through a learnable Kalman Filtering statistical theory-based approach.
 Extensive empirical studies are conducted on three benchmark STG forecasting datasets, demonstrating the performance superiority and computational efficiency of STG-Mamba.
 It not only surpasses existing state-of-the-art methods in terms of STG forecasting performance, but also effectively alleviate the computational bottleneck of large-scale graph networks in reducing the computational cost of FLOPs and test inference time.
+The implementation code is available at: https://github.com/LincanLi98/STG-Mamba.
 1 Introduction
@@
 (2023) further demonstrates the power of GNNs in predicting air quality, emphasizing the importance of group-aware mechanisms in enhancing STG forecasting performance.
-In the realm of Social Recommendation, Huang et al.
-(2021) introduces a coupled GNN framework that incorporates inter-dependent knowledge graphs across both items and users.
-Wang et al.
-(2020a) explores the use of disentangled representations within a heterogeneous graph attention network to refine recommendation systems.
+In the realm of Social Recommendation, Zhang et al.
+(2023) introduces an Adaptive Disentangled Transformer (ADT) framework that optimizes the disentanglement of attention heads in different layers of Transformer architecture.
+Wei et al.
+(2023) explores the improved efficiency of multimedia recommendation by introducing modal-specific embeddings and a lightweight self-attention.
 Although Transformer-based methods demonstrate notable improvements in STG relationship learning Chen et al.
@@
 SSSM is a variant of the State Space Models, which originated from Control Science and Engineering fields Lee et al.
-(1994); Friedland (2012) with strong theoretical foundation.
+(1994); Friedland (2012).
 State Space Models provide a professional framework to describe a physical system’s dynamics state evolution through sets of input, output, and state variables related by first-order differential or difference equations.
 The approach allows for a compact way to model and analyze systems with multiple inputs and outputs (MIMO) Aoki (2013).
-As a pioneer deep learning-based SSSM, Mamba is proposed to address data-centric context learning for sequential tasks.
-Different from attention score calculation, Mamba construct selective state space models for the research objects, which encode contexts by using recurrent hidden states.
-The selection mechanism within Mamba allows it to control which part of the input data should be flowed into hidden states, as the significant context features that will influence the next-step embedding update.
 STG learning can be treated as a complex process of understanding and forecasting the evolution of STG networks over time, which highly resembles the state space transition process where each state encapsulates the current configuration of a system and transitions represent changes over time.
@@
 A simple yet elegant way is employed to extend selective state space models to handle STG data.
-Specifically, we formulate the framework as Encoder-Decoder architecture, with $N$ stacked GS3B as basic modules and the proposed Kalman Filtering Graph Neural Network (KFGN) within it, to facilitate efficient Seq2Seq modeling.
-  • A novel Graph Selective State Space Block (GS3B) is proposed for STG network pioneering integration with selective state space models, which performs input-dependent edge construction and dynamic node feature selection.
-GS3B is capable of capturing the sophisticated heterogenous dependencies among STG network, and accurately modeling the dynamic evolution of STG system over time, making it highly fitted for STG learning.
-  • We introduce Kalman Filtering Graph Neural Network (KFGN) as the specialized adaptive spatiotemporal graph generation and upgrading method, which smoothly fits in within the SSSM-based context, ensuring the synchronous updates between graph structure and STG systems’s current state.
+Specifically, we formulate the framework following the Stacked Residual Encoder Fashion, with $N$ stacked Graph Selective State Space Block (GS3B) as basic modules and the proposed Kalman Filtering Graph Neural Network (KFGN) within it, to facilitate efficient Seq2Seq modeling.
+  • A novel Spatial-Temporal Selective State Space Module (ST-S3M) is proposed for STG network pioneering integration with selective state space models, which performs input-dependent adaptive spatial-temporal graph feature selection.
+ST-S3M is capable of capturing the sophisticated heterogenous dependencies among STG network, and accurately modeling the dynamic evolution of STG system over time, making it highly fitted for STG learning.
+  • We introduce KFGN as the specialized adaptive spatiotemporal graph generation and upgrading method, which smoothly fits in within the SSSM-based context.
+In KFGN, the Dynamic Filter Linear serves as dynamic GNN module which generating input-specific graph structures.
+The KF-Upgrading mechanism models STG inputs from different temporal granularities as parallel streams, and the output embeddings are integrated through the statistical theory learning of Kalman Filtering for optimization.
   • Extensive evaluations are carried on three open-sourced benchmark STG dataset.
@@
 STG is a kind of non-Euclidean structured data, which extensively existed in daily-life, business, and society.
-For a given graph network, if the node individuals have certain spatial connectivities at any given time step $\tau$, and the node features and/or graph structures are dynamically evolving over temporal axis, it could be defined as an STG network.
+For a given graph network, if the node individuals have certain spatial connectivity at any given time step $\tau$, and the node features and/or graph structures are dynamically evolving over temporal axis, it could be defined as an STG network.
 Formally, we use $\mathbb{G}^{ST}=(V^{ST},E^{ST},A^{ST})$ to represent a given STG network.
@@
 For STG System, this framework is adapted to accommodate spatial dependencies and temporal evolution, structuring the system with nodes representing spatial entities, edges denoting spatial interactions, and state variables evolving over time to capture the dynamics of STG data.
-State Space Models.
-State Space Models (SSMs) are a foundational concept in control science and system identification, offering a comprehensive framework to describe the dynamic evolutions of systems through mathematical theories.
-SSMs articulate a system’s behavior by defining a set of equations that represent the system’s state evolution over time, incorporating both control inputs and outputs.
-Formally, SSMs are expressed through state equations $x_{t+1}=f(x_{t},u_{t})$ and observation equations $y_{t}=g(x_{t},u_{t})$, where $x_{t}$ denotes the state at time $t$, $u_{t}$ represents the control input, and $y_{t}$ is the observed output.
-The state transition function $f(\cdot)$ describes how the current state $x_{t}$ evolves to the next state $x_{t+1}$ under the influence of the control input $u_{t}$.
-$f(\cdot)$ captures the dynamics of the system, accounting for how internal and external factors modify the system’s state over time.
-The observation function $g(\cdot)$ is used for mapping the current state $x_{t}$ and control input $u_{t}$ to the observed output $y_{t}$.
-This function defines how the underlying state of the system is reflected in the measurements or observations collected.
-As a variant of SSMs, Selective State Space Model emphasizes how to build the selection mechanism upon SSMs, which is highly resemble to the core idea of attention mechanism, making it a competitive counterpart of Transformer architecture.
+State Space Models (SSMs).
+Deep Learning-based SSMs are a lately invented class of sequential models that are closely related to RNN architectures and classical state space models.
+They are characterized by a particular continuous system model that maps a multi-dimensional input sequence to a corresponding output sequence through an implicit latent state representation.
+SSMs are defined by four parameters $(\mathbb{A},\mathbb{B},\mathbb{C},\mathbb{D})$ that specify how the input (control signal) and the current state determine the next state and output.
+This framework allows for efficient sequential modeling by enabling both linear and non-linear computations.
+As a variant of SSMs, SSSMs emphasize how to build the selection mechanism upon SSMs, which is highly resemble to the core idea of attention mechanism, making it a competitive counterpart of Transformer architecture.
 STG Forecasting based on SSSMs.
-When employing Selective State Space Models (SSSMs) for spatial-temporal graph forecasting, the problem can be formulated as dynamically identifying and utilizing relevant portions of historical spatiotemporal data and graph structures to predict the future states of a STG system.
+Employing SSSMs for spatial-temporal graph forecasting, the problem can be formulated as dynamically identifying and utilizing relevant portions of historical spatiotemporal data and graph structures to predict the future states of a STG system.
 Given spatiotemporal graph $\mathbb{G}^{ST}=(V^{ST},E^{ST},A^{ST})$ and historical spatial-temporal sequential data $\rm{X}_{t-p+1:t}=\{X_{t-p+1},X_{t-p+2},...,X_{t}\}$, the objective is to utilize SSSMs to forecast future STG system’s states $\rm{\hat{X}_{t+1:t+k}}=\{\hat{X}_{t+1},...,\hat{X}_{t+k}\}$.
@@
 In this section, we start with introducing the model architecture of STG-Mamba and the formulation of STG Selective State Space Models.
-Following that, we elaborate the key contributions of this work, including the Kalman Filtering Graph Neural Networks (Section 3.2) which specifically designed for SSSMs to enable dynamic adaptive graph learning; and the Graph Selective State Space Blocks (GS3B) which acts as the basic modules for STG-Mamba (Section 3.3).
+Following that, we elaborate the key contributions of this work, including the Kalman Filtering Graph Neural Networks (Section 3.2) which specifically designed for SSSMs to enable adaptive statistical-based graph learning under contexts with noisy data and uncertainty; and the Spatial-Temporal Selective State Space Module (ST-S3M) which serves as the state space selection mechanism for spatial-temporal graph learning.
 Finally, we describe the computational efficiency of STG-Mamba in Section 3.4.
@@
 Figure 1: The comprehensive illustration of STG-Mamba’s Model Architecture.
-Modern State Space Models all rely on a classical continuous system that maps a one-dimensional input function or sequence, denoted as $x(t)\in\mathbb{R}$, through intermediate implicit states $h(t)\in\mathbb{R}^{N}$ to an output $y(t)\in\mathbb{R}$.
-The aforementioned process can be represented as a linear Ordinary Differential Equation (ODE) as in Eq.
-2:
-$\begin{array}[]{lr}h^{\prime}(t)=\mathbb{A}h_{t}+\mathbb{B}x(t)\\ y(t)=\mathbb{C}h(t)\end{array}$ (2)
-where $\mathbb{A}\in\mathbb{R}^{N\times N}$ represents the State Matrix, while $\mathbb{B}\in\mathbb{R}^{N\times 1},\mathbb{C}\in\mathbb{R}^{N\times 1}$ denote the Input Matrix and Output Matrix, respectively.
-Obtaining the output sequence $y(t)$ at a given time-step $t$ necessitates determining $h(t)$ first, which poses great analytical challenges.
-As such, modern selective state space models address this challenge by converting the continuous system into a discretized version, enhancing the suitability for deep learning scenarios while boosting computational efficiency.
-A critical step in the discretization process involves the introduction of a time-scale factor $\Delta$, leading to the discrete representation of matrices $\mathbb{A}$ and $\mathbb{B}$ as $\overline{\mathbb{A}}$ and $\overline{\mathbb{B}}$, respectively, following a pre-determined discretization method, commonly the zero-order hold (ZOH) approach.
-$\begin{array}[]{lr}\overline{\mathbb{A}}=\exp(\Delta\mathbb{A})\\ \overline{\mathbb{B}}=(\Delta\mathbb{A})^{-1}(\exp(\Delta A)-\mathbb{I})\cdot\Delta\mathbb{B}\end{array}$ (3)
-After discretization, SSM-based methods can be computed in two ways: linear recurrent or global convolution, defined as in Eq. 4 and Eq. 5, respectively:
-$\begin{array}[]{lr}h^{\prime}(t)=\overline{\mathbb{A}}h_{t}+\overline{\mathbb{B}}x(t)\\ y(t)=\mathbb{C}h(t)\end{array}$ (4)
-$\begin{array}[]{lr}\overline{\mathbb{K}}=(\mathbb{C\overline{B}},\mathbb{C\overline{AB}},...,\mathbb{C\overline{A}^{L-1}\overline{B}})\\ y=x*\overline{\mathbb{K}}\end{array}$ (5)
-where $\overline{K}\in\mathbb{R}^{L}$ denotes a structured convolutional kernel, and $L$ denotes the length of the input sequence $x$.
+Modern selective state space models, such as Mamba, build upon classical continuous systems by incorporating deep learning techniques and discretization methods suitable for complex time-series data.
+Mamba specifically maps a one-dimensional input sequence $x(t)\in\mathbb{R}$ through intermediate implicit states $h(t)\in\mathbb{R}^{N}$, culminating in an output $y(t)\in\mathbb{R}$.
+The fundamental process is described by a modified set of state space equations that include discretization adjustments to accommodate deep learning frameworks:
+$\begin{array}[]{lr}h^{\prime}(t)=\overline{\mathbb{A}}h(t)+\overline{\mathbb{B}}x(t)\\ y(t)=\mathbb{C}h(t)\end{array}$ (2)
+where $\overline{\mathbb{A}}\in\mathbb{R}^{N\times N}$ and $\overline{\mathbb{B}}\in\mathbb{R}^{N\times 1}$ are discretized representations of the continuous state transition and input matrices, respectively.
+$\mathbb{C}\in\mathbb{R}^{1\times N}$ is the output matrices.
+The discretization process involves transformation techniques that convert $\mathbb{A}$ and $\mathbb{B}$ from their continuous forms to computational efficient formats.
+These transformations are determined by a selective scan algorithm that utilizes a convolutional approach to model interactions across temporal dimension:
+$\begin{array}[]{lr}\overline{\mathbb{A}}=\exp(\Delta\mathbb{A})\\ \overline{\mathbb{B}}=\Delta\mathbb{B}\cdot(\exp(\Delta\mathbb{A})-\mathbb{I})\end{array}$ (3)
+Mamba extends classic state space model functionalities by incorporating a selective mechanism that dynamically adapts to the input sequence characteristics.
+This adaptability is achieved through a novel architecture that selectively applies convolution kernels across the sequence, optimizing the model’s response to spatiotemporal variations:
+$\begin{array}[]{lr}\overline{\mathbb{K}}=(\mathbb{C}\overline{\mathbb{B}},\mathbb{C}\overline{\mathbb{A}\mathbb{B}},\dots,\mathbb{C}\overline{\mathbb{A}^{L-1}\mathbb{B}})\end{array}$ (4)
+where $\overline{\mathbb{K}}$ denotes a structured convolutional kernel, capturing dependencies across spatial and temporal dimensions, and $L$ represents the length of the input sequence $x$.
+This selective convolutional approach significantly enhances the model’s capability to forecast spatiotemporal data by focusing computation on critical features of the data sequence, thus improving both accuracy and efficiency in predictions.
 Figure 1 illustrates the comprehensive architecture of the proposed STG-Mamba.
-Specifically, we formulate the overall architecture following the Encoder-Decoder fashion for efficient sequential data modeling and prediction.
-In STG-Mamba, we leverage the Graph Selective State Space Block (GS3B) as the basic module, and repeat it $N$ times for Encoder and Decoder part.
-GS3B consists of several operations, including Layer Norm, Linear Layer, Kalman Filter Graph Neural Networks (KFGN), Graph State Space Layer (GSS-Layer), and Residual Connections.
-We specifically design KFGN and GSS-Layer for STG learning task build upon selective state space models.
+Specifically, we formulate the overall architecture following the Residual Encoder fashion for efficient sequential data modeling and prediction.
+In STG-Mamba, we leverage the Graph Selective State Space Block (GS3B) as the basic Encoder module, and repeat it $N$ times.
+GS3B consists of several networks and operations, including Layer Norm, Kalman Filter Graph Neural Networks (KFGN) which consists of DynamicFilter-GNN followed by Kalman Filter Upgrading (KF-Upgrading), Spatial-Temporal Selective State Space Module (ST-S3M) which consists of Linear & Split, Conv1D, SiLU, State Space Selection Mechanism (SSSM), Element-Wise Concatenation, as well as th Residual Connections for every GS3B Encoder layer.
 3.2 Kalman Filtering Graph Neural Networks
-Due to the dynamic, heterogenous, and non-stationary characteristics of STG data, the spatiotemporal dependencies of an STG system also evolve over time.
-When reflecting such graph dynamics in GNN formulation, it means the time-variant dynamic graph adjacency matrix.
-In existing literature, researchers usually employ various kinds of Attention-based methods to construct dynamic graph adjacency matrices, which may not be feasible in the context of state space model-based STG learning.
-On the other hand, Kalman Filtering (KF), as a well-known application of SSMs, has been extensively used in sequential data filtering and forecasting.
-In light of that, we propose a novel approach which employs KF-based method to generate dynamic-evolving graphs, which delicately captures the system’s state-space transition for a given STG network.
-The method is termed as Kalman Filtering Graph Neural Networks (KFGN).
-To generate graph adjacency matrix $\mathbb{\alpha_{t}}$ at each time step using KFGN, the modeling procedure involves two steps as shown in the following:
-Step 1: State Equation Calculation: Let $\mathbb{\alpha_{t}}$ denotes the graph adjacency matrix at time $t$, which indicates the current state of the STG system.
-The state equation is defined as $\mathbb{\alpha_{t}}=F_{t-1}\mathbb{\alpha_{t-1}}+W_{t-1}$, where $F_{t-1}$ is the state transition matrix that reflects changes in the graph adjacency matrix from $t-1$ to $t$, and $W_{t-1}$ represents the noise of the KF transition procedure.
-Step 2: Observation Equation Calculation: The observed STG feature can be modeled as $X_{t}^{{}^{\prime}}=H_{t}\mathbb{\alpha_{t}}+V_{t}$, where $X_{t}^{{}^{\prime}}$ is the observation of STG feature at time $t$, $H_{t}$ is the observation matrix, and $V_{t}$ represents observation noise.
-Iteratively applying these two equations through Kalman Filter enables dynamic estimation of the graph adjacency matrix $\alpha_{t}$ at each moment, adaptively capturing the evolving dynamics of the STG network.
-In practice, we employ a Neural Network-based method called Deep Kalman Filter (DKF) to automatically learn the state transition matrix $F_{t-1}$ and observation matrix $H_{t}$.
-Specifically, we adopt Gated Recurrent Unit (GRU) network as the architecture of Deep Kalman Filter.
-The GRU-based DKF is employed to learn the non-linear mapping for state transition and observation generation.
-The state transition matrix $F_{t-1}$ is modeled by GRU inner process to capture the non-linear relationships between states, predicting the state for the next moment.
-Similarly, the observation matrix $H_{t}$ is predicted under the current state by GRU.
-This process enables the system to automatically learn the dynamics of state transitions and observation generation from data without manually settings.
-After receiving the $\alpha_{t}$, the next step of KFGN is the graph convolution operation, which is defined as:
-$h_{t}^{kfgn}=(\omega_{gc}\odot\alpha_{t})x_{t}^{{}^{\prime}}$ (6)
-where $\odot$ is the Hadamard product, and $x_{t}^{{}^{\prime}}$ denotes the input feature of KFGN at time $t$. $\omega_{gc}$ is the learnable weight matrix of KFGN.
-3.3 GS3B: Graph Selective State Space Blocks
-Figure 2: The selection mechanism of STG-Mamba that facilitates input-dependent feature selection and focusing.
-Inspired from Mamba, the GS3B is the major component of STG-Mamba, as depicted in Figure 1.
-After being processed by Layer Normalization, the input is divided into two sub-branches.
-In the first branch, the input STG embedding passed through a Linear Layer, followed by the proposed KFGN network and an activation function, before entering the GSS-Layer for selective feature learning.
-After that, the learned features are normalized by the Layer Norm.
-For the second branch, the input STG embedding undergoes a Linear Layer and an activation function.
-Subsequently, the two branches are fused together by an element-wise dot product operation.
-Finally, the merged embedding passes through a linear layer, and the outcome is combined with a residual connection to form the GS3B block’s output.
-In this research, Sigmoid Linear Unit (SiLU) is employed as the default activation function.
-We further comprehend the detailed process of the STG-Mamba’s Selection Mechanism in Figure 2.
-The GSS-Layer is a graph selective state space module specifically designed for STG data.
-GSS-Layer incorporates a selective mechanism which is capable of adjusting the SSM model’s parameters according to the input features.
-The characteristics enable the model to distinguish and retain key information which contributes significantly in STG learning tasks, while filtering out the irrelevant information.
-The pseudo-code of GSS-Layer (graph selective state space algorithm) is elaborated in Algorithm 1.
-Algorithm 1 Pseudo Code of the Proposed Graph Selective State Space Algorithm
-Input: $X\in\mathbb{R}^{batch\times L\times dim}$, the historical ground-truth STG data, where $batch$ is batch size of model training, $L$ is the historical time window length, $dim$ is the feature dimension of STG data;
-$\mathbb{\alpha_{t}}$, the graph adjacency matrix of the current time step calculated by Deep Kalman Filter Network.
-Params: $\mathbb{A}$ is the network param: nn.Parameter; $\mathbb{D}$ is the network param: nn.Parameter;
-Operator: $Linear(\cdot)$, the linear projection layer;
-Output:$y$, the processed feature with shape $\mathbb{R}^{batch\times L\times dim}$;
-1:  $\Delta,\mathbb{B},\mathbb{C}=Linear(X),Linear(X),Linear(X)$
-2:  $\overline{\mathbb{A}}=\exp(\Delta\mathbb{A})$
-3:  $\overline{\mathbb{B}}=(\Delta\mathbb{A})^{-1}(\exp(\Delta\mathbb{A})-I)\cdot\Delta B$
-4:  $h_{t}=(\overline{\mathbb{A}}\odot\mathbb{\alpha_{t}})h_{t-1}+\overline{\mathbb{B}}x_{t}$
-5:  $y_{t}=\mathbb{C}h_{t}+\mathbb{D}x_{t}$
-6:  $y=[y_{1},y_{2},...,y_{t},...,y_{L}]$.
-7:  return $y$
+The motivation for employing Kalman Filtering-based methods in constructing GNNs stems from the need to enhance the reliability and accuracy of spatial-temporal forecasting.
+STG big data (such as traffic sensor records/weather station records/urban metro system In/Out records) often contains inherent biases and noises that existing methods typically overlook.
+By integrating Kalman Filtering-based optimization and upgrading, our KFGN effectively addresses these inaccuracies by dynamically weighing the reliability of data streams from different temporal granularity, optimizing the fusion of these data streams based on the estimated variances.
+This approach not only corrects dataset inherent errors but also significantly improves the model’s capacity to capture the complex inter-dependencies within STG patterns.
+As depicted in Figure 2, the KFGN pipeline consists of two key steps.
+In the first step, embeddings generated from the model inputs of different temporal granularities (i.e., recent steps/periodic steps/trend steps) are sent into a DynamicFilter-GNN module for processing.
+Then, in the second step, the outputs of DynamicFilter-GNN module are integrated and optimized through the Kalman Filtering Upgrading module.
+In the initial stage of DynamicFilter-GNN, a learnable Base Filter parameter matrix of size $\mathbb{R}^{\text{in\_fea}\times\text{in\_fea}}$ is defined.
+It’s employed to transform the graph adjacency matrix, thereby dynamically adjusting the connection degrees between nodes.
+Following that, we uniformly initialize the weights and biases, with an initialization standard deviation stdv calculated as the reciprocal of the square root of the number of input features: $\text{stdv}=\frac{1}{\sqrt{\text{in\_fea}}}$.
+This uniform initialization is critical for ensuring the module starts from a neutral point.
+Figure 2: Kalman Filtering Graph Neural Networks.
+Subsequently, we transformer the Base Filter via a linear transformation layer, and combine it with the original adjacency matrix $A_{\text{ini}}^{c_{i}}$, resulting in a dynamically adjusted adjacency matrix $A_{\text{dyn}}^{c_{i}}$.
+Let $c_{i,i=\{1,2,3\}}\in\{r,p,q\}$ represents the input data from different temporal granularities, (i.e., $c_{1}=r$ denotes historical recent, $c_{2}=p$ denotes historical period, $c_{3}=q$ denotes historical trend).
+Leveraging $A_{\text{dyn}}^{c_{i}}$, input embedding $h_{in}^{c_{i}}$, weights $W_{DF}^{c_{i}}$ and bias $b_{DF}^{c_{i}}$, the input embedding undergoes a graph convolution:
+$h_{DF}^{c_{i}}=h_{in}^{c_{i}}\cdot(A_{\text{dyn}}^{c_{i}}\cdot W_{DF}^{c_{i}})+b_{DF}^{c_{i}}$ (5)
+This design enables the model to dynamically adjust the strengths of connections between nodes within the graph based on the learned adjustments.
+Following DynamicFilter-GNN, the next step is Kalman Filtering Upgrading.
+Since the input data are subjected to the same dataset but within different temporal granularities, they are assumed to follow Gaussian Distributions:
+$y_{q}(x;\mu_{q},\sigma_{q})\triangleq\frac{\exp^{-\frac{(x-\mu_{q})^{2}}{2\sigma_{q}^{2}}}}{\sqrt{2\pi\sigma_{q}^{2}}};y_{p}(x;\mu_{p},\sigma_{p})\triangleq\frac{\exp^{-\frac{(x-\mu_{p})^{2}}{2\sigma_{p}^{2}}}}{\sqrt{2\pi\sigma_{p}^{2}}};y_{r}(x;\mu_{r},\sigma_{r})\triangleq\frac{\exp^{-\frac{(x-\mu_{r})^{2}}{2\sigma_{r}^{2}}}}{\sqrt{2\pi\sigma_{r}^{2}}};$ (6)
+where subscript $q=c_{3},p=c_{2},r=c_{1}$ denote historical trend/period/recent dataset, respectively.
+We employ Kalman Filtering method to derive accurate information from the multi-granularity observation sets.
+Specifically, their individual probability distribution functions can be integrated together by multiplication:
+$y_{fuse}(x;\mu_{q},\sigma_{q},\mu_{p},\sigma_{p},\mu_{r},\sigma_{r})$ $=\frac{\exp^{-\frac{(x-\mu_{q})^{2}}{2\sigma_{q}^{2}}}}{\sqrt{2\pi\sigma_{q}^{2}}}\times\frac{\exp^{-\frac{(x-\mu_{p})^{2}}{2\sigma_{p}^{2}}}}{\sqrt{2\pi\sigma_{p}^{2}}}\times\frac{\exp^{-\frac{(x-\mu_{r})^{2}}{2\sigma_{r}^{2}}}}{\sqrt{2\pi\sigma_{r}^{2}}}$
+$=\frac{1}{(2\pi)^{3/2}\sqrt{\sigma_{q}^{2}\sigma_{p}^{2}\sigma_{r}^{2}}}\exp^{-\left(\frac{(x-\mu_{q})^{2}}{2\sigma_{q}^{2}}+\frac{(x-\mu_{p})^{2}}{2\sigma_{p}^{2}}+\frac{(x-\mu_{r})^{2}}{2\sigma_{r}^{2}}\right)}$ (7)
+By reorganizing Eq. 7 into a simplified version, we have:
+$y_{fuse}(x;\mu_{fuse},\sigma_{fuse})=\frac{1}{\sqrt{2\pi\sigma_{fuse}^{2}}}\exp^{-\frac{(x-\mu_{fuse})^{2}}{2\sigma_{fuse}^{2}}}$ (8)
+where $\mu_{fuse}=\frac{\mu_{q}/\sigma_{q}^{2}+\mu_{p}/\sigma_{p}^{2}+\mu_{r}/\sigma_{r}^{2}}{1/\sigma_{q}^{2}+1/\sigma_{p}^{2}+1/\sigma_{r}^{2}}$ and $\sigma_{fuse}^{2}=\frac{1}{1/\sigma_{q}^{2}+1/\sigma_{p}^{2}+1/\sigma_{r}^{2}}$.
+To simplify the representation of $\mu_{fuse}$ and $\sigma_{fuse}^{2}$, we hereby introduce parameters $\omega_{q}=\frac{1}{\sigma_{q}^{2}},\omega_{p}=\frac{1}{\sigma_{p}^{2}},\omega_{r}=\frac{1}{\sigma_{r}^{2}}$.
+Then, $\mu_{fuse}$ and $\sigma_{fuse}^{2}$ can be re-written as:
+$\mu_{fuse}=\frac{\mu_{q}\omega_{q}+\mu_{p}\omega_{p}+\mu_{r}\omega_{r}}{\omega_{q}+\omega_{p}+\omega_{r}};\\ \sigma_{fuse}^{2}=\frac{1}{\omega_{q}+\omega_{p}+\omega_{r}}$ (9)
+This means that observations from different branches can be effectively integrated using weighted sum, and the weight is a combination of variances:
+$\mu_{fuse}$ $=$ $\mu_{q}(\frac{\omega_{q}}{\omega_{q}+\omega_{p}+\omega_{r}})+\mu_{p}(\frac{\omega_{p}}{\omega_{q}+\omega_{p}+\omega_{r}})+\mu_{r}(\frac{\omega_{r}}{\omega_{q}+\omega_{p}+\omega_{r}})$ (10)
+$\Downarrow$
+$y_{fuse}$ $=$ $y_{q}(\frac{\omega_{q}}{\omega_{q}+\omega_{p}+\omega_{r}})+y_{p}(\frac{\omega_{p}}{\omega_{q}+\omega_{p}+\omega_{r}})+y_{r}(\frac{\omega_{r}}{\omega_{q}+\omega_{p}+\omega_{r}})$
+Directly calculating the variance of an observation set requires the complete data, which is computationally expensive.
+Therefore, we estimate the variance by computing the variance distribution of each training samples:
+$\mathbb{E}[\sigma_{\{q,p,r\}}^{2}]=\frac{1}{N_{m}}\sum_{i}\frac{(S_{i}-\overline{S})^{2}}{L}$ (11)
+where $L$ is the length of each sample sequence, and $N_{m}$ is the number of data samples.
+$S_{i}$ denotes the $i$-th observation, and $\overline{S}$ denotes the average value of all observed samples.
+To further improve the integration and upgrading, we add two learnable weight params $\epsilon,\varphi$ to balance different observation branches.
+To this end, based on the formulation of $y_{fuse}$ in Eq.
+10, the output of Kalman Filtering Upgrading module is:
+$\tilde{y}_{fuse}=\frac{\epsilon\cdot(\hat{y}_{q}\omega_{q})+\varphi\cdot(\hat{y}_{p}\omega_{p})+\hat{y}_{r}\omega_{r}}{\omega_{q}+\omega_{p}+\omega_{r}}$ (12)
+where $\omega_{q}=\frac{1}{\sigma_{q}^{2}},\omega_{p}=\frac{1}{\sigma_{p}^{2}},\omega_{r}=\frac{1}{\sigma_{r}^{2}}$, $\epsilon$ and $\varphi$ are trainable weights.
+Finally, for the convenience of neural network training and ensure the scalability, we remove the constant denominator $(\omega_{q}+\omega_{p}+\omega_{r})$ in Eq.
+12.
+As such, the final output of Kalman Filtering Upgrading is defined as:
+$\hat{y}_{fuse}=\epsilon\cdot(\hat{y}_{q}\omega_{q})+\varphi\cdot(\hat{y}_{p}\omega_{p})+\hat{y}_{r}\omega_{r}$ (13)
+Here, $\hat{y}_{fuse}$ is the final output of KF-Upgrading module.
+Note that our approach is an optimized version of the classic Kalman Filtering theory, which is specifically designed for deep learning-based methodology, ensuring both computationally efficiency, dynamic hierarchical STG feature fusion, and improved accuracy.
+3.3 ST-S3M: Spatial-Temporal Selective State Space Module
+Figure 3: Illustration of the spatial-temporal state space selection mechanism that facilitates input-dependent feature selection and focusing.
+ST-S3M plays the significant role of adaptive STG feature selection, serving as the counterpart of Attention mechanism.
+The architecture of ST-S3M is shown in Figure 1.
+We also give an vivid illustration of the STG feature selection procedure in Figure 4.
+After being processed by KFGN module for dynamic spatiotemporal dependency modeling and statistical-based integration & upgrading.
+The generated embedding $\hat{h}_{fuse}$ is sent to ST-S3M for input-specific dynamic feature selection.
+In ST-S3M, the input STG embedding passes through a Linear Layer, followed by a splitting operation.
+Let $b$ denotes batch size, $l$ denotes sequence length, $d_{model}$ be the feature dimension, $d_{inner}$ be the model inner feature dimension, we have:
+$h_{\text{main-res}}$ $=W_{in}\hat{h}_{fuse}+b_{in}$
+$(h_{\text{main}},res)$ $=\text{split}(h_{\text{main-res}})$ (14)
+where $\hat{h}_{fuse}\in\mathbb{R}^{b\times l\times d_{model}}$, $W_{in}\in\mathbb{R}^{2d_{inner}\times d_{model}}$, $b_{in}\in\mathbb{R}^{2d_{inner}}$, $h_{\text{main-res}}\in\mathbb{R}^{b\times l\times 2d_{inner}}$, $h_{\text{main}},res\in\mathbb{R}^{b\times l\times d_{inner}}$
+Then, $h_{\text{main}}$ flows into a 1D convolution layer, followed by SiLU activation function:
+$h_{\text{main}}^{{}^{\prime}}=\text{Conv1D}(h_{\text{main}})$
+$h_{\text{main}}^{{}^{\prime\prime}}=\text{SiLU}(h_{\text{main}}^{{}^{\prime}})$ (15)
+where $h_{\text{main}}^{{}^{\prime\prime}}\in\mathbb{R}^{b\times l\times d_{inner}}$. The output of SiLU is sent to state space selection mechanism for processing:
+$h_{\text{sssm}}=\text{SSSM}(h_{main}^{{}^{\prime\prime}})$ (16)
+where $h_{\text{sssm}}\in\mathbb{R}^{b\times l\times d_{inner}}$.
+On the same time, the residual part $res$ is also processed by SiLU activation, and finally we employ element-wise dot product to fuse the main STG embedding and its residual: $h_{\text{sssm}}\odot\text{SiLU}(res)$.
+The fused result is transformed through a Linear projection:
+$h_{out}=W_{out}(h_{\text{sssm}}\odot\text{SiLU}(res))+b_{out}$ (17)
+where $h_{out}\in\mathbb{R}^{b\times l\times d_{model}}$ is the final output of ST-S3M. $W_{out}\in\mathbb{R}^{d_{model}\times d_{inner}}$, $b_{out}\in\mathbb{R}^{d_{model}}$.
+The State Space Selection Mechanism in ST-S3M plays a pivotal role in adaptive spatiotemporal feature selection.
+We detail the parameter computation and updating process of the State Space Selection Mechanism in Algorithm 1.
+Further, the pseudo-code for the employed Selective Scan method is presented in Algorithm 2.
+Algorithm 1 State Space Selection Mechanism Parameter Calculation and Upgrading.
+Input: $x\in\mathbb{R}^{b\times l\times d_{\text{model}}}$, learnable model params $A_{log}\in\mathbb{R}^{d_{\text{in}}\times n}$ and $D\in\mathbb{R}^{d_{\text{in}}}$;
+Output: $y\in\mathbb{R}^{b\times l\times d_{\text{inner}}}$;
+1:  Compute input-independent params $A\in\mathbb{R}^{d_{\text{in}}\times n}$ and $D\in\mathbb{R}^{d_{\text{in}}}$:
+2:  $A=-\exp(A_{log})$, $D=D$;
+3:  Process input $x$ with Linear Layer $Input\_proj$:
+4:  $x_{\text{dbl}}=Input\_proj(x)$, $x_{\text{dbl}}\in\mathbb{R}^{b\times l\times(dt_{\text{rank}}+2n)}$;
+5:  Dividing $x_{\text{dbl}}$ in order to get the input-dependent params $\Delta,B,C$, where $\Delta\in\mathbb{R}^{b\times l\times dt_{\text{rank}}}$, $B,C\in\mathbb{R}^{b\times l\times n}$;
+6:  Use Linear Layer $dt\_proj$ and $SoftPlus$ activation to adjust $\Delta$:
+7:  $\Delta^{*}=SoftPlus(dt\_proj(\Delta))$, where $\Delta^{*}\in\mathbb{R}^{b\times l\times d_{\text{in}}}$;
+8:  Employ the $SelectiveScan$ function to proceed with $x,\Delta^{*},A,B,C,D$:
+9:  $y=SelectiveScan(x,\Delta^{*},A,B,C,D)$;
+10:  Return the Output $y$.
+Algorithm 2 Selective Scan Algorithm
+Input: $u,\Delta^{*},A,B,C,D$;
+Output: $z$;
+1:  Obtain the dim of $u$ as $(b,l,d_{\text{in}})$ and the second dim of $A$ as $n$;
+2:  Discretize continuous parameters $A$ and $B$:
+3:  $\Delta A\leftarrow\exp(\text{einsum}(\Delta^{*},A))$, where $\Delta A\in\mathbb{R}^{b\times l\times d_{\text{in}}\times n}$;
+4:  $\Delta B_{u}\leftarrow\text{einsum}(\Delta^{*},B,u)$, where $\Delta B_{u}\in\mathbb{R}^{b\times l\times d_{\text{in}}\times n}$;
+5:  Initialize state $x$ as zeroes with dimension $\mathbb{R}^{b\times d_{\text{in}}\times n}$;
+6:  for $i$ from $1$ to $l$ do
+7:     $x\leftarrow\Delta A[:,i]\times x+\Delta B_{u}[:,i]$;
+8:     $z\leftarrow\text{einsum}(x,C[:,i,:])$;
+9:     Append $z$ to the output list $z_{s}$;
+10:  end for
+11:  $z\leftarrow\text{stack}(z_{s})$, where $z\in\mathbb{R}^{b\times l\times d_{\text{in}}}$;
+12:  Adding the direct gain $D$:
+13:  $z\leftarrow z+u\times D$
+14:  Return $z$.
 3.4 Computational Efficiency of STG-Mamba
-GS3B’s selection mechanism is introduced in Algorithm 1, which outlines the graph selective state space algorithm for efficiently processing STG data.
-It starts with initializing linear projections of the input STG data to obtain parameters $\Delta,\mathbb{B},\mathbb{C}$, which are used to compute the modified state transition matrix $\overline{\mathbb{A}}$ and input matrix $\overline{\mathbb{B}}$.
-These matrices incorporate the graph adjacency matrix $\mathbb{\alpha_{t}}$, updated by a Deep Kalman Filter Network, to calculate the hidden state $h_{t}$ for each time step.
-Finally, the output $y_{t}$ is generated by applying linear transformations to the hidden state and the input data, aggregating the results over the historical time window to form the output feature $y$.
-This algorithm effectively integrates dynamic graph structure information into SSSM, enhancing its capability to model sophisticated STG dynamics.
 Leveraging a data-specific selection process, the increasing of sequence length $L$ or graph network node scale $N$ by $k$-fold can results in huge expansion of model parameters for Transformer-based methods.
@@
 (2022), and weather station records across mainland China (KnowAir) Wang et al.
-(2020b).
+(2020).
 Acquired from the Caltrans Performance Measurement System (PeMS), PeMS04 records the average traffic speed from a total number of 307 sensors distributed on California Highway system.
@@
 Baseline Methods. For fair comparison, we consider both benchmark spatiotemporal graph neural network (STGNN)-based methods, and Transformer-based methods for STG learning.
-STGNN-based Methods:
-(I) STGCN Yu et al.
-(2018): STGCN propose to capture the complex spatial-temporal dependencies within traffic network by graph neural networks, and design a fully convolutional approach for fast training.
-(II) STSGCN Song et al. (2020): STSGCN presents a novel model which efficiently handling complex localized spatial-temporal correlations and data heterogeneity through a synchronous mechanism.
-(III) STG-NCDE Choi et al.
-(2022): STG-NCDE leverages the concept of Neural Controlled Differential Equations (NCDEs) for spatial-temporal processing within a unified framework, achieving superior accuracy across six benchmark datasets.
-(IV) DDGCRN Weng et al.
-(2023): DDGCRN dynamically captures the spatial-temporal features by separating the normal and abnormal signals and employing a novel graph convolutional recurrent network approach.
-Attention (Transformer)-based Methods:
-(V) ASTGCN Guo et al.
-(2019): ASTGCN leverages an attention-based spatial-temporal graph convolutional network to accurately predict STG data by dynamically capturing complex spatiotemporal correlations and patterns.
-(VI) ASTGNN Guo et al.
-(2021): Attention based Spatial-Temporal Graph Neural Network incorporates a unique trend-aware self-attention and dynamic graph convolution to address temporal dynamics, spatial correlations, and the inherent periodicity and heterogeneity of STG data.
-(VII) PDFormer Jiang et al.
-(2023): PDFormer introduces a tailored Transformer model, adept at capturing dynamic spatial dependencies and addressing the time delay in traffic condition propagation with its unique spatial self-attention and traffic delay-aware feature transformation.
-(VIII) STAEformer Liu et al.
-(2023): STAEformer introduces a novel approach that enhances vanilla Transformer with a spatiotemporal adaptive embedding component, enabling it to capture complex STG patterns.
-(IX) MultiSPANS Zou et al.
-(2024): MultiSPANS leverages multi-filter convolution modules, Transformers, and structural entropy optimization to effectively model complex multi-range dependencies in STG forecasting.
+STGNN-based Methods: (I) STGCN Yu et al. (2018), (II) STSGCN Song et al. (2020), (III) STG-NCDE Choi et al. (2022), (IV) DDGCRN Weng et al. (2023).
+Attention (Transformer)-based Methods: (V) ASTGCN Guo et al.
+(2019), (VI) ASTGNN Guo et al.
+(2021), (VII) PDFormer Jiang et al.
+(2023), (VIII) STAEformer Liu et al.
+(2023), (IX) MultiSPANS Zou et al.
+(2024).
 4.2 Implementation Settings
@@
 Before starting model training session, all the data samples are normalized into range $[0,1]$ with MinMax normalization method.
-We set the batch size as 32 and employ AdamW as the Optimizer.
-The learning rate is initialized as $1e^{-3}$, with a weight decay of $5e^{-2}$.
+We set the batch size as 48 and employ AdamW as the Optimizer.
+The learning rate is initialized as $1e^{-4}$, with a weight decay of $1e^{-2}$.
 CosineAnnealingLR Loshchilov and Hutter (2016) is adopted as the learning rate scheduler, with a maximum of 50 iterations, and a minimum learning rate $1e^{-5}$.
-Model training epochs is set as 300.
-The number of Encoder/Decoder Layer is set as $N=4$ and $N^{\prime}=4$, respectively.
-We use the historical ground-truth data from the past 12 steps to forecast one future time step.
+Model training epochs is set as 100.
+The number of GS3B Encoder Layer is set as $N=4$.
+We use the historical ground-truth data from the past 12 steps (for historical recent/period/trend scales) to forecast the future 12 time steps.
 Mean Squared Error (MSE) is employed as the Loss Function for optimization.
@@
 4.3 Results Evaluation and Comparison
-4.3.1 Experimental Results Comparison on Three STG Forecasting Tasks
 A comprehensive evaluation between STG-Mamba and the proposed baseline methods is conducted, and we report the RMSE/MAE/MAPE performance results in Table 2.
@@
 STG-Mamba 29.53 18.09 12.12 29.26 18.38 11.67 18.05 11.73 10.66
-4.3.2 Computational Efficiency Evaluation
-STG-Mamba offers substantial improvement in computational efficiency in addition to performance gain.
-In order to quantatively evaluate computational efficiency, we select the widely recognized Floating Point Operations (FLOPs) and Inference Time as criteria.
-For both of the two criteria, a smaller value indicates better model performance.
-STG-Mamba is compare with other Transformer-based benchmark methods: ASTGNN, PDFormer, and STAEformer, on the same testing set.
-Table 3 showcases the inference time comparison on PeMS04 and KnowAir testing set.
-It can be seen that ASTGNN is relatively redundancy, with the lowest inference time among the compared Transformer-based methods.
-STG-Mamba gains 18.09% and 20.54% performance increase in inference time than the 2nd best performance model (STAEformer).
-We further compare the computational FLOPs increment with accordance to the scale complexity of STG network.
-Specifically, we use the number of node individuals to represent the scale of STG network.
-In the experiments, we select different number of nodes from PeMS04 and KnowAir.
-We select $[50,100,150,200,250,300]$ and $[30,60,90,120,150,180]$ node individuals for PeMS04 and KnowAir dataset, respectively.
-STG-Mamba is compared with the benchmark STAEformer in every settings.
-Figure 4 visualizes the FLOPs evaluation under different STG scale complexity.
-We can easily find that STG-Mamba enjoys a linear increment in FLOPs, while STAEformer shows quadratic increment in FLOPs.
-At the initial small-scale setting, the FLOPs for STG-Mamba/STAEformer are quite similar.
-However, with the increasing of network complexity and scale, the difference becomes more significant.
-At the 300 node setting of PeMS04, the FLOPs for STG-Mamba and STAEformer is 36.85G and 92.49G (2.50 times more), respectively.
-At the 180 node setting of KnowAir, the FLOPs for STG-Mamba and STAEformer is 57.08G and 118.13G (2.07 times more), respectively.
-Benefiting from the linear computational increment brought by modern selective state space models, STG-Mamba and alike models demonstrate stronger capability and computational efficiency for modeling large-scale STG data, which holds promise for addressing the computational overhead in future large-scale STG learning.
-Table 3: Test inference time comparison with benchmark Transformer-based approaches.
-Model Inference Time on PeMS04 Inference Time on KnowAir
-ASTGNN 4.85 s 5.31s
-PDFormer 3.93 s 4.28s
-STAEformer 3.76 s 4.09s
-STG-Mamba 3.08 s 3.25s
-performance $\uparrow$ +18.08% +20.54%
-Figure 4: (Computational Efficiency Eval) FLOPs comparison between STG-Mamba and STAEformer on PeMS04/KnowAir dataset under different STG node number settings.
 4.4 Robustness Analysis
@@
   2. 2.
-    STG-Mamba w/o KFGN: We replace the proposed Kalman Filtering Graph Neural Networks (KFGN) within STG-Mamba with the standard Graph Convolution Network (GCN).
+    STG-Mamba w/o KF-Upgrading: We replace the KF-Upgrading module in KFGN with the simple sum average operation to fuse the three temporal branches.
   3. 3.
-    STG-Mamba w/o GSS-Layer: The GSS-Layer with proposed Graph Selective State Space Algorithm is replaced by the standard selective state space module as introduced in Gu and Dao (2023).
+    STG-Mamba w/o Dynamic Filter Linear: We replace the proposed Dynamic Filter Linear in KFGN with the basic static graph convolution.
   4. 4.
-    Transformer w KFGN: The GS3B block is reformulated as Transformer block (similar to ASTGNN Guo et al. (2021)), which consists of attention mechanism and the proposed KFGN.
-As illustrated in Figure 5, STG-Mamba w/o KFGN shows slight worse performance than the full model, demonstrating the necessity and suitability of employing Kalman filtering-based dynamic GNN for SSSM-based methods.
-Different from the mainstream methods which employ various attention mechanisms to upgrade STG network’s adjacency matrix, our KFGN is more consistent with the SSSM context, and at the same time keep the linear algorithm complexity.
-STG-Mamba w/o GSS-Layer performs even worse than w/o KFGN.
-This is because in our design, GSS-Layer will take the generated STG adjacency matrix $\alpha_{t}$ at this timestamp as input for selective state space algorithm’s better upgradation.
-Replacing GSS-Layer with the standard selective state space layer leads to a less suitable version for STG learning.
-Finally, when we change from Mamba-based architecture to Transformer-based architecture, there is a substantial performance drop among all the three tasks.
-This discovery proves the effectiveness of modern SSSM over attention-based designs.
+    STG-Mamba w/o SSSM: The State Space Selection Mechansim (SSSM) is removed from the ST-S3M module.
+  5. 5.
+    STG-Mamba w/o ST-S3M: The whole ST-S3M module is removed from GS3B Encoder.
+As illustrated in Figure 5, STG-Mamba w/o KF-Upgrading shows worse performance than the full model, demonstrating the necessity and suitability of employing statistical learning-based GNN state space upgrading and optimization for SSSM-based methods.
+The decrease of performance due to the absence of Dynamic Filter Linear demonstrate its effectiveness, and the necessity of designing a suitable adaptive GNN for STG feature learning.
+Furthermore, removing the State Space Selection Mechansim in ST-S3M results in a substantial degradation of model capability, proving the feasibility of using SSSM as the substitute for Attention mechanism.
+Finally, the removal of ST-S3M makes the STG-Mamba degrade to a plain GNN-based model, resulting in the lowest performance.
 5 Conclusions
-In this work, we first propose a novel spatial-temporal graph (STG) learning model named STG-Mamba that leverages modern selective state space models (SSSM) for accurate and efficient STG forecasting tasks.
-STG-Mamba is formulated as Encoder-Decoder structure, with the Graph Selective State Space Block (GS3B) as basic module.
-GS3B’s selection method facilitates input-dependent edge construction and node feature selection, successfully integrates STG network with SSSMs.
-Through the Kalman Filtering Graph Neural Networks (KFGN), the STG network achieves a smooth and dynamic evolution process of spatiotemporal graph adjacency matrix upgrading, which align with the whole graph selective state space modeling process.
-Compared with Attention-based methods, our STG-Mamba is able to achieve linear-time complexity and substantial reduction in FLOPs and test inference time.
-We conduct extensive empirical studies on three open-sourced STG forecasting datasets from urban traffic, metro In/Out flow, and weather records.
-Experimental results demonstrate the consistent superiority of STG-Mamba over other benchmark methods in both forecasting performance and computational efficiency.
+In this work, we for the first time introduce deep learning-based selective state space models (SSSM) for spatial-temporal graph learning tasks.
+We propose STG-Mamba that leverages modern SSSM for accurate and efficient STG forecasting.
+STG-Mamba is formulated following the stacked Residual Encoder architecture, with the ST-S3M module facilitating input-dependent graph evolution and node feature selection, successfully integrates STG network with SSSMs.
+Through the Kalman Filtering Graph Neural Networks (KFGN), the learned STG embeddings achieve a smooth optimization & upgrading based on statistical learning, aligning with the whole STG selective state space modeling process.
+Compared with Attention-based methods, STG-Mamba is able to achieve linear-time complexity and substantial reduction in FLOPs and test inference time.
+We conduct extensive empirical studies on three open-sourced STG datasets from urban traffic, metro In/Out flow, and weather records.
+Experimental results demonstrate the consistent superiority of STG-Mamba over other benchmark methods.
 We believe that STG-Mamba presents a brand new promising approach to general STG learning fields, offering competitive model performance and input-dependent context-aware feature selection under an affordable computational cost.

### v4

@@
 Spatial-temporal graph data is a kind of Non-Euclidean data which widely exists in our daily life, such as urban traffic network, metro system InFlow/OutFlow, social networks, regional energy load, weather observations, etc.
-Owning to the strong ability of capturing dynamic features across both temporal and spatial dimension, attention-based methods, such as Spatial-Temporal Transformer and its variants have recently been broadly implemented in spatial-temporal graph (STG) learning tasks.
-In the domain of STG traffic forecasting, seminal works have emerged that leverage these methods to enhance prediction accuracy.
-ASTGCN Guo et al.
-(2019) introduces an innovative approach combining spatial-attention & temporal attention mechanisms with graph neural networks.
-ADCT-Net Kong et al.
-(2024) presents to incorporate a dual-graphic cross-fusion Transformer for dynamic traffic forecasting.
-Huo et al.
-(2023) integrates hierarchical graph convolutional networks with Transformer networks to capture sophisticated STG relationships among traffic data.
-For weather forecasting, Liang et al.
-(2023) exemplifies how Transformer-based models can be utilized for accurate air quality forecasting across vast geographical areas.
-Chen et al.
-(2023) further demonstrates the power of GNNs in predicting air quality, emphasizing the importance of group-aware mechanisms in enhancing STG forecasting performance.
-In the realm of Social Recommendation, Zhang et al.
-(2023) introduces an Adaptive Disentangled Transformer (ADT) framework that optimizes the disentanglement of attention heads in different layers of Transformer architecture.
-Wei et al.
-(2023) explores the improved efficiency of multimedia recommendation by introducing modal-specific embeddings and a lightweight self-attention.
-Although Transformer-based methods demonstrate notable improvements in STG relationship learning Chen et al.
-(2022); Cong et al.
-(2021), their modeling ability on large-scale STG networks and Long-Term Sequential Forecasting (LTSF) tasks is greatly hindered by the quadratic computational complexity $O(n^{2})$ provided by its core component–attention mechanism Zhou et al.
-(2021, 2022); Rampášek et al.
-(2022).
-Moreover, attention mechanisms usually approximate the full-length historical data, or encode all the context features, which may not be necessary for modeling the long-term dependencies of STG learning tasks.
-As such, the research world has been actively looking for counterparts of Attention-based models.
+Owning to the dynamic, heterogenous, and non-stationary nature of STG data, accurate and efficient STG forecasting has long been a challenging task.
 Lately, with the popularization of Mamba Gu and Dao (2023); Wang et al.
 (2024); Liu et al.
-(2024), Selective State Space Models (SSSM) have aroused considerable interest among researchers from computer vision and Natural Language Processing.
+(2024), Modern Selective State Space Models (SSSM) have aroused considerable interest among researchers from computer vision and Natural Language Processing.
 SSSM is a variant of the State Space Models, which originated from Control Science and Engineering fields Lee et al.
@@
 Deep learning-based SSSMs bring new horizons to STG learning tasks.
-However, great challenges pose in accurately and effectively adopting Mamba architectures for STG modeling, which is characterized as heterogeneous, dynamic, and non-stationary.
+However, great challenges pose in accurately and effectively adopting SSSMs architectures for STG modeling.
 Motivated by the excellent long-term contextual modeling ability and low computational overhead of SSSMs, we propose Spatial-Temporal Graph Mamba (STG-Mamba) as the first deep learning-based SSSM for effective data-centric STG learning.
-The main contributions of this work are summarized in the following:
+The main contributions of this work are:
   • We make the first attempt to adapting SSSMs for STG learning tasks.
-A simple yet elegant way is employed to extend selective state space models to handle STG data.
-Specifically, we formulate the framework following the Stacked Residual Encoder Fashion, with $N$ stacked Graph Selective State Space Block (GS3B) as basic modules and the proposed Kalman Filtering Graph Neural Network (KFGN) within it, to facilitate efficient Seq2Seq modeling.
+A simple yet elegant way is employed to extend SSSMs to handle STG data.
+Specifically, we formulate the framework following the Stacked Residual Encoder Fashion, which takes $N$ stacked Graph Selective State Space Block (GS3B) as basic module, with the proposed Kalman Filtering Graph Neural Network (KFGN), the Spatial-Temporal Selective State Space Module (ST-S3M), and a simultaneous STG Feed-Forward connection to serialize and coordinate different internal modules.
   • A novel Spatial-Temporal Selective State Space Module (ST-S3M) is proposed for STG network pioneering integration with selective state space models, which performs input-dependent adaptive spatial-temporal graph feature selection.
-ST-S3M is capable of capturing the sophisticated heterogenous dependencies among STG network, and accurately modeling the dynamic evolution of STG system over time, making it highly fitted for STG learning.
+The Graph Selective Scan algorithm within ST-S3M simultaneously receives graph information from KFGN through the Feed-Forward connection, the Feed-Forward graph information is then employed to assist more effective updating of the state transition matrix and control matrix.
   • We introduce KFGN as the specialized adaptive spatiotemporal graph generation and upgrading method, which smoothly fits in within the SSSM-based context.
-In KFGN, the Dynamic Filter Linear serves as dynamic GNN module which generating input-specific graph structures.
+In KFGN, DynamicFilter-GNN serves as the module which generating input-specific dynamic graph structures.
 The KF-Upgrading mechanism models STG inputs from different temporal granularities as parallel streams, and the output embeddings are integrated through the statistical theory learning of Kalman Filtering for optimization.
@@
 2 Preliminaries
-Spatial-Temporal Graph (STG).
-STG is a kind of non-Euclidean structured data, which extensively existed in daily-life, business, and society.
-For a given graph network, if the node individuals have certain spatial connectivity at any given time step $\tau$, and the node features and/or graph structures are dynamically evolving over temporal axis, it could be defined as an STG network.
-Formally, we use $\mathbb{G}^{ST}=(V^{ST},E^{ST},A^{ST})$ to represent a given STG network.
-Here, $V^{ST}$ denotes the set of graph vertices (nodes) within the network, $E^{ST}$ denotes the edges of the graph network.
-$A^{ST}$ is the graph adjacency matrix, which indicates the calculated proximity between node individuals.
 Spatial-Temporal Graph System (STG system).
-We for the first time define a network which consists of STG data as a Spatial-Temporal Graph System (STG system).
-In the theory foundation, a system is defined as the representation of a physical process, comprising state variables that describe the system’s current condition, input variables influencing the system’s state, and output variables reflecting the system’s response.
+We for the first time define a network which consists of STG data as a Spatial-Temporal Graph System.
+In theory foundation, a system is defined as the representation of a physical process, comprising state variables that describe the system’s current condition, input variables influencing the system’s state, and output variables reflecting the system’s response.
 For STG System, this framework is adapted to accommodate spatial dependencies and temporal evolution, structuring the system with nodes representing spatial entities, edges denoting spatial interactions, and state variables evolving over time to capture the dynamics of STG data.
@@
 3 Methodology
-In this section, we start with introducing the model architecture of STG-Mamba and the formulation of STG Selective State Space Models.
-Following that, we elaborate the key contributions of this work, including the Kalman Filtering Graph Neural Networks (Section 3.2) which specifically designed for SSSMs to enable adaptive statistical-based graph learning under contexts with noisy data and uncertainty; and the Spatial-Temporal Selective State Space Module (ST-S3M) which serves as the state space selection mechanism for spatial-temporal graph learning.
-Finally, we describe the computational efficiency of STG-Mamba in Section 3.4.
-3.1 Formulation of STG Selective State Space Models
-SSSMs like structured state space sequential models (S4) and Mamba, initially designed for sequential data, are adapted for Spatial-Temporal Graph (STG) data modeling by leveraging their structure to handle spatiotemporal dependencies dynamically.
-The adaptation involves discretizing the continuous system described by linear Ordinary Differential Equations (ODEs) into a format amenable to deep neural networks modeling and optimization, enhancing computational efficiency while capturing the evolving dynamics of the STG networks.
+In this section, we start with introducing the model architecture of STG-Mamba and the formulation of STG Selective State Space Models in Section 3.1 and Appendix D as well.
+Following that, we elaborate the key contributions of this work, including the Kalman Filtering Graph Neural Networks (Section 3.2) which specifically designed for SSSMs to enable adaptive statistical-based graph learning under contexts with noisy data and uncertainty; and the Spatial-Temporal Selective State Space Module (ST-S3M) which serves as the state space selection mechanism for spatial-temporal graph learning (Section 3.3).
+Finally, a detailed computational efficiency analysis of STG-Mamba in presented in Appendix B because of limited space.
+3.1 Architecture of STG Selective State Space Models
 Figure 1: The comprehensive illustration of STG-Mamba’s Model Architecture.
-Modern selective state space models, such as Mamba, build upon classical continuous systems by incorporating deep learning techniques and discretization methods suitable for complex time-series data.
-Mamba specifically maps a one-dimensional input sequence $x(t)\in\mathbb{R}$ through intermediate implicit states $h(t)\in\mathbb{R}^{N}$, culminating in an output $y(t)\in\mathbb{R}$.
-The fundamental process is described by a modified set of state space equations that include discretization adjustments to accommodate deep learning frameworks:
-$\begin{array}[]{lr}h^{\prime}(t)=\overline{\mathbb{A}}h(t)+\overline{\mathbb{B}}x(t)\\ y(t)=\mathbb{C}h(t)\end{array}$ (2)
-where $\overline{\mathbb{A}}\in\mathbb{R}^{N\times N}$ and $\overline{\mathbb{B}}\in\mathbb{R}^{N\times 1}$ are discretized representations of the continuous state transition and input matrices, respectively.
-$\mathbb{C}\in\mathbb{R}^{1\times N}$ is the output matrices.
-The discretization process involves transformation techniques that convert $\mathbb{A}$ and $\mathbb{B}$ from their continuous forms to computational efficient formats.
-These transformations are determined by a selective scan algorithm that utilizes a convolutional approach to model interactions across temporal dimension:
-$\begin{array}[]{lr}\overline{\mathbb{A}}=\exp(\Delta\mathbb{A})\\ \overline{\mathbb{B}}=\Delta\mathbb{B}\cdot(\exp(\Delta\mathbb{A})-\mathbb{I})\end{array}$ (3)
-Mamba extends classic state space model functionalities by incorporating a selective mechanism that dynamically adapts to the input sequence characteristics.
-This adaptability is achieved through a novel architecture that selectively applies convolution kernels across the sequence, optimizing the model’s response to spatiotemporal variations:
-$\begin{array}[]{lr}\overline{\mathbb{K}}=(\mathbb{C}\overline{\mathbb{B}},\mathbb{C}\overline{\mathbb{A}\mathbb{B}},\dots,\mathbb{C}\overline{\mathbb{A}^{L-1}\mathbb{B}})\end{array}$ (4)
-where $\overline{\mathbb{K}}$ denotes a structured convolutional kernel, capturing dependencies across spatial and temporal dimensions, and $L$ represents the length of the input sequence $x$.
-This selective convolutional approach significantly enhances the model’s capability to forecast spatiotemporal data by focusing computation on critical features of the data sequence, thus improving both accuracy and efficiency in predictions.
 Figure 1 illustrates the comprehensive architecture of the proposed STG-Mamba.
@@
 In STG-Mamba, we leverage the Graph Selective State Space Block (GS3B) as the basic Encoder module, and repeat it $N$ times.
-GS3B consists of several networks and operations, including Layer Norm, Kalman Filter Graph Neural Networks (KFGN) which consists of DynamicFilter-GNN followed by Kalman Filter Upgrading (KF-Upgrading), Spatial-Temporal Selective State Space Module (ST-S3M) which consists of Linear & Split, Conv1D, SiLU, State Space Selection Mechanism (SSSM), Element-Wise Concatenation, as well as th Residual Connections for every GS3B Encoder layer.
+GS3B consists of several networks and operations, including Layer Norm, Kalman Filter Graph Neural Networks (KFGN) which consists of DynamicFilter-GNN followed by Kalman Filter Upgrading (KF-Upgrading), Spatial-Temporal Selective State Space Module (ST-S3M) which consists of Linear & Split, Conv1D, SiLU, Grapg State Space Selection Mechanism (GSSSM), Element-Wise Concatenation, Graph Information Feed-Forward, as well as the Residual Connection.
+The Graph Information Feed-Forward structure here is specifically designed for coordinating different modules’ information transmission and updating, ensuring the latest STG information is available for every module.
 3.2 Kalman Filtering Graph Neural Networks
 The motivation for employing Kalman Filtering-based methods in constructing GNNs stems from the need to enhance the reliability and accuracy of spatial-temporal forecasting.
-STG big data (such as traffic sensor records/weather station records/urban metro system In/Out records) often contains inherent biases and noises that existing methods typically overlook.
-By integrating Kalman Filtering-based optimization and upgrading, our KFGN effectively addresses these inaccuracies by dynamically weighing the reliability of data streams from different temporal granularity, optimizing the fusion of these data streams based on the estimated variances.
+STG big data (i.e., traffic sensor records, weather station records, etc) often contains inherent biases and noises that existing methods typically overlook.
+By integrating Kalman Filtering-based optimization and upgrading, KFGN effectively addresses these inaccuracies by dynamically weighing the reliability of data streams from different temporal granularity, optimizing the fusion of these data streams based on the estimated variances.
 This approach not only corrects dataset inherent errors but also significantly improves the model’s capacity to capture the complex inter-dependencies within STG patterns.
@@
 To this end, based on the formulation of $y_{fuse}$ in Eq.
-10, the output of Kalman Filtering Upgrading module is:
+7, the output of Kalman Filtering Upgrading module is:
 $\tilde{y}_{fuse}=\frac{\epsilon\cdot(\hat{y}_{q}\omega_{q})+\varphi\cdot(\hat{y}_{p}\omega_{p})+\hat{y}_{r}\omega_{r}}{\omega_{q}+\omega_{p}+\omega_{r}}$ (9)
@@
 Finally, for the convenience of neural network training and ensure the scalability, we remove the constant denominator $(\omega_{q}+\omega_{p}+\omega_{r})$ in Eq.
-12.
+9.
 As such, the final output of Kalman Filtering Upgrading is defined as:
@@
 3.3 ST-S3M: Spatial-Temporal Selective State Space Module
-Figure 3: Illustration of the spatial-temporal state space selection mechanism that facilitates input-dependent feature selection and focusing.
 ST-S3M plays the significant role of adaptive STG feature selection, serving as the counterpart of Attention mechanism.
 The architecture of ST-S3M is shown in Figure 1.
-We also give an vivid illustration of the STG feature selection procedure in Figure 4.
 After being processed by KFGN module for dynamic spatiotemporal dependency modeling and statistical-based integration & upgrading.
@@
 Then, $h_{\text{main}}$ flows into a 1D convolution layer, followed by SiLU activation function:
-$h_{\text{main}}^{{}^{\prime}}=\text{Conv1D}(h_{\text{main}})$
-$h_{\text{main}}^{{}^{\prime\prime}}=\text{SiLU}(h_{\text{main}}^{{}^{\prime}})$ (15)
-where $h_{\text{main}}^{{}^{\prime\prime}}\in\mathbb{R}^{b\times l\times d_{inner}}$. The output of SiLU is sent to state space selection mechanism for processing:
-$h_{\text{sssm}}=\text{SSSM}(h_{main}^{{}^{\prime\prime}})$ (16)
+$h_{\text{main}}^{{}^{\prime}}=\text{SiLU}(\text{Conv1D}(h_{\text{main}}))$ (12)
+where $h_{\text{main}}^{{}^{\prime}}\in\mathbb{R}^{b\times l\times d_{inner}}$. The output of SiLU is sent to graph state space selection mechanism:
+$h_{\text{sssm}}=\text{GSSSM}(h_{main}^{{}^{\prime}})$ (13)
 where $h_{\text{sssm}}\in\mathbb{R}^{b\times l\times d_{inner}}$.
-On the same time, the residual part $res$ is also processed by SiLU activation, and finally we employ element-wise dot product to fuse the main STG embedding and its residual: $h_{\text{sssm}}\odot\text{SiLU}(res)$.
+Meanwhile, the residual part $res$ is also processed by SiLU activation, and finally we employ element-wise dot product to fuse the main STG embedding and its residual: $h_{\text{sssm}}\odot\text{SiLU}(res)$.
 The fused result is transformed through a Linear projection:
@@
 where $h_{out}\in\mathbb{R}^{b\times l\times d_{model}}$ is the final output of ST-S3M. $W_{out}\in\mathbb{R}^{d_{model}\times d_{inner}}$, $b_{out}\in\mathbb{R}^{d_{model}}$.
-The State Space Selection Mechanism in ST-S3M plays a pivotal role in adaptive spatiotemporal feature selection.
-We detail the parameter computation and updating process of the State Space Selection Mechanism in Algorithm 1.
-Further, the pseudo-code for the employed Selective Scan method is presented in Algorithm 2.
-Algorithm 1 State Space Selection Mechanism Parameter Calculation and Upgrading.
-Input: $x\in\mathbb{R}^{b\times l\times d_{\text{model}}}$, learnable model params $A_{log}\in\mathbb{R}^{d_{\text{in}}\times n}$ and $D\in\mathbb{R}^{d_{\text{in}}}$;
-Output: $y\in\mathbb{R}^{b\times l\times d_{\text{inner}}}$;
-1:  Compute input-independent params $A\in\mathbb{R}^{d_{\text{in}}\times n}$ and $D\in\mathbb{R}^{d_{\text{in}}}$:
-2:  $A=-\exp(A_{log})$, $D=D$;
-3:  Process input $x$ with Linear Layer $Input\_proj$:
-4:  $x_{\text{dbl}}=Input\_proj(x)$, $x_{\text{dbl}}\in\mathbb{R}^{b\times l\times(dt_{\text{rank}}+2n)}$;
-5:  Dividing $x_{\text{dbl}}$ in order to get the input-dependent params $\Delta,B,C$, where $\Delta\in\mathbb{R}^{b\times l\times dt_{\text{rank}}}$, $B,C\in\mathbb{R}^{b\times l\times n}$;
-6:  Use Linear Layer $dt\_proj$ and $SoftPlus$ activation to adjust $\Delta$:
-7:  $\Delta^{*}=SoftPlus(dt\_proj(\Delta))$, where $\Delta^{*}\in\mathbb{R}^{b\times l\times d_{\text{in}}}$;
-8:  Employ the $SelectiveScan$ function to proceed with $x,\Delta^{*},A,B,C,D$:
-9:  $y=SelectiveScan(x,\Delta^{*},A,B,C,D)$;
-10:  Return the Output $y$.
-Algorithm 2 Selective Scan Algorithm
-Input: $u,\Delta^{*},A,B,C,D$;
-Output: $z$;
-1:  Obtain the dim of $u$ as $(b,l,d_{\text{in}})$ and the second dim of $A$ as $n$;
-2:  Discretize continuous parameters $A$ and $B$:
-3:  $\Delta A\leftarrow\exp(\text{einsum}(\Delta^{*},A))$, where $\Delta A\in\mathbb{R}^{b\times l\times d_{\text{in}}\times n}$;
-4:  $\Delta B_{u}\leftarrow\text{einsum}(\Delta^{*},B,u)$, where $\Delta B_{u}\in\mathbb{R}^{b\times l\times d_{\text{in}}\times n}$;
-5:  Initialize state $x$ as zeroes with dimension $\mathbb{R}^{b\times d_{\text{in}}\times n}$;
-6:  for $i$ from $1$ to $l$ do
-7:     $x\leftarrow\Delta A[:,i]\times x+\Delta B_{u}[:,i]$;
-8:     $z\leftarrow\text{einsum}(x,C[:,i,:])$;
-9:     Append $z$ to the output list $z_{s}$;
-10:  end for
-11:  $z\leftarrow\text{stack}(z_{s})$, where $z\in\mathbb{R}^{b\times l\times d_{\text{in}}}$;
-12:  Adding the direct gain $D$:
-13:  $z\leftarrow z+u\times D$
-14:  Return $z$.
-3.4 Computational Efficiency of STG-Mamba
-Leveraging a data-specific selection process, the increasing of sequence length $L$ or graph network node scale $N$ by $k$-fold can results in huge expansion of model parameters for Transformer-based methods.
-STG-Mamba introduces a sophisticated algorithm that is conscious of hardware capabilities, exploiting the layered structure of GPU memory to minimize the ensuing cost.
-In detail, for a batch size of $B$, STG feature dimension of $d^{\prime}$, and network node number of $N$, STG-Mamba reads the $O(BLd^{\prime}+Nd^{\prime})$ input data of $\mathbb{A},\mathbb{B},\mathbb{C},\Delta$ from High Bandwidth Memory (HBM), processes the intermediate stages sized $O(BLd^{\prime}N)$ in Static Random-Access Memory (SRAM), and dispatches the ultimate output, sized $O(BLd^{\prime})$, back to HBM.
-In contrast, if a Transformer-based model is employed, the HBM will read $O(BL^{2}d^{\prime})$ input data, and the computational size remains $O(BL^{2}d^{\prime})$ in SRAM.
-In common practices we often have $BL^{2}d\gg BLd^{\prime}+Nd^{\prime}$, which means SSSM-based approach can effectively decrease the computational overhead in large-scale parallel processing, such as neural network training.
-Additionally, by not retaining intermediate states, it also diminishes memory usage, since these states are recalculated during the gradient assessment in reverse phase.
-SSSM’s GPU-optimized design ensures STG-Mamba operates with a linear time complexity $O(L)$ relative to the length of the input sequence, marking a substantial speed advantage over traditional transformers’ dense attention computations, which exhibit quadratic time complexity $O(L^{2})$.
+The GSSSM in ST-S3M plays a main role in adaptive spatiotemporal feature selection.
+We detail the parameter computation and update process of the Graph State Space Selection Mechanism in Algorithm 2.
+Graph Selective Scan Algorithm (Algorithm 1) which receives graph information feed-forward is the most significant step in the state space selection process.
+We detail it in the following.
+Graph Selective Scan Algorithm.
+The algorithm is a further extension of the basic selective scan, which integrates dynamic graph information generated by KFGN into state-space selection & update procedure, enhancing Mamba’s capability in capturing STG dependencies.
+The key steps and modifications are as follows:
+We start with obtaining the dimension of the input tensor $u\in\mathbb{R}^{(b,l,d_{in})}$, where $b$ is the batch size, $l$ denotes the sequence length, $d_{in}$ is the input features dim; and the second dim of $\mathbb{A}$ is denoted as $n$.
+[…]
+Following that, we Discretize the continuous parameters $\mathbb{A}$ and $\mathbb{B}$ as:
+[…]
+[…]
+where einsum denotes the Einstein Summation Convention. $\text{delta}\mathbb{A}$ is the updated state transition matrix, $\text{delta}\mathbb{B}_{u}$ is the updated control matrix.
+Then, Iterative State Update is executed on state $x\in\mathbb{R}^{b\times d_{in}\times n}$. For time step from $i$ to $l$, we have:
+$x$ $\leftarrow\text{delta}\mathbb{A}[:,i]\times x+\text{delta}\mathbb{B}_{u}[:,i]$ (16)
+$z$ $\leftarrow\text{einsum}(x,\mathbb{C}[:,i,:])$
+where $z$ is the current output computed by einstein summation convention.
+$z$ is then appended to the output list $z_{s}$.
+The list of outputs $z_{s}$ is stacked to form the output tensor $z$, and finally we add the direct gain $\mathbb{D}$ to the final output as: $z\leftarrow z+u\times\mathbb{D}$.
+The Graph Selective Scan algorithm offers several advantages including: enhanced spatiotemporal dependency modeling, adaptability to changing graph structures, and improved forecasting accuracy, making it particularly suitable for spatiotemporal graph learning tasks.
 4 Experiments
@@
 Datasets.
-For experimental evaluation, we meticulously select three real-world STG datasets from California road network speed records (PeMS04) Song et al.
-(2020), Hangzhou urban metro system entry/exit records (HZMetro) Liu et al.
-(2022), and weather station records across mainland China (KnowAir) Wang et al.
-(2020).
-Acquired from the Caltrans Performance Measurement System (PeMS), PeMS04 records the average traffic speed from a total number of 307 sensors distributed on California Highway system.
-The time range of PeMS04 is 1st/Jan/2018 - 28th/Feb/2018, with 5-minute time interval.
-HZMetro was created with the In/Out transaction records of 80 stations of Hangzhou Metro system from 1st Jan 2019 to 25th Jan 2019, which takes 15-minute as data collection time interval.
-KnowAir documents weather observations in every 3-hour, encompassing a dataset that spans from 2015 to 2018.
-Within the KnowAir dataset, there are three subsets.
-In this study, we exclusively use subset No.
-3, covering the period from 1/Sep/2016 to 31/Jan/2017.
+We meticulously select three real-world STG datasets from California road network speed records, Hangzhou metro system entry/exit records, and weather station records across mainland China.
+Namely PeMS04 Song et al.
+(2020), HZMetro Liu et al.
+(2022), and KnowAir Wang et al.
+(2020), respectively.
 The detailed dataset statistics and descriptions are summarized in Table 1.
@@
 Mean Squared Error (MSE) is employed as the Loss Function for optimization.
-And finally, we employ RMSE, MAE, MAPE as the evaluation metrics.
 All experiments are conducted under one NVIDIA A100 80GB GPU for neural computing acceleration.
+4.3 Evaluation Metrics
+We employ RMSE/MAE/MAPE/$std_{\text{MAE}}$ as the deterministic evaluation metrics, all the indicators are smaller the better; $\rm{\Delta_{\text{VAR}}}$ (Change in Variance Accounted For) and $\rm{R^{2}}$ (Coefficient of Determination) as statistic-based evaluation metric, where $\rm{\Delta_{\text{VAR}}}$ is the smaller the better, $\rm{R^{2}}$ is the larger the better.
+The reported results in Experiment section are the average value of 10 runs.
+The detailed indicator calculation methods are provided in Appendix B because of limited space.
 4.4 Results Evaluation and Comparison
-A comprehensive evaluation between STG-Mamba and the proposed baseline methods is conducted, and we report the RMSE/MAE/MAPE performance results in Table 2.
-Our STG-Mamba consistently outperforms other benchmark methods on the three tasks, except for the MAPE criteria of PeMS04 (Flow).
-In PeMS04 (Flow) forecasting task, STAEformer shows slightly better MAPE results than STG-Mamba.
+A comprehensive evaluation between STG-Mamba and the proposed baseline methods is conducted, and we report the RMSE/MAE/MAPE/$std_{\text{MAE}}$ performance results in Table 2.
+STG-Mamba consistently outperforms other benchmark methods, except for the MAPE criteria of PeMS04 (Flow).
 It can also be easily observed that Transformer-based methods which integrate GNNs inside the architecture always gain superior performance than sole GNN-based methods.
@@
 Within the GNN-based methods, STG-NCDE and DDGCRN stand out for their competitive performance.
-STG-NCDE models continuous-time dynamics within STG networks through neural controlled differential equations (NCDE), offering a more precise representation of spatiotemporal feature changes over time compared to other discrete GNN-based methods.
-Figure 4 presents a direct performance visualization comparison among STG-Mamba and other superior Transformer-based baselines (ASTGNN/PDFormer/STAEformer) on PeMS04, HZMetro, and KnowAir dataset, respectively.
-We visualize the predictions over 72 consecutive time steps using the average results of multiple 12-step forecast settings.
-For each dataset, four node individuals are randomly selected (i.e., 4 random traffic sensors of PeMS04, 4 random metro stations of HZMetro, and 4 random weather stations of KnowAir).
-The results indicate that STG-Mamba consistently yields the prediction curve closest to the ground truth waveform, featuring more accurate STG sequence peak values and fewer fluctuations.
-Figure 4: STG forecasting results visualization and comparison among STG-Mamba and other superior Transformer-based baselines.
-As a new contender poised to challenge Transformer architecture, Mamba introduces an innovative approach by replacing the foundational Attention mechanism with modern selective state space model (SSSM).
-SSSM diverges from Attention by utilizing a continuous-time model that captures the dynamic evolution of spatial-temporal dependencies more naturally and efficiently.
-Unlike Attention, which computes weights across all pairs of input and output positions and can be computationally intensive, Mamba’s selective state space model directly captures the temporal evolution of features, enabling it to scale more gracefully with the sequence length and complexity of spatial-temporal data.
-Furthermore, we facilitate STG-Mamba with Deep Kalman Filtering-based adaptive spatiotemporal graph evolution, and the elaborately designed Graph Selective State Space Block (GS3B), making it inherit the advantages of modern SSSM, and highly suitable for STG learning tasks.
-Table 2: Performance Evaluation and Comparison with Benchmark Baselines on PeMS04, HZMetro, KnowAir Datasets.
+STG-NCDE models continuous-time dynamics through neural controlled differential equations (NCDE), offering a more precise representation of spatiotemporal feature changes over time compared to other discrete GNN-based methods.
+As a new contender poised to challenge Transformer architecture, the selective state space mechanism of Mamba diverges from Attention by utilizing a continuous-time model that captures the dynamic evolution of spatial-temporal dependencies more naturally and efficiently.
+Unlike Attention, SSSM directly captures the temporal evolution of features, enabling it to scale more gracefully with the sequence length and complexity of STG data.
+Furthermore, we facilitate STG-Mamba with Kalman Filtering-based adaptive STG evolution, the elaborately designed graph selective scan algorithm, and the Feed-Forward connection, making it inherit the advantages of modern SSSM, and highly suitable for STG learning.
+Table 2: Performance Eval and Comparison with Baselines on PeMS04/HZMetro/KnowAir Dataset.
 Model PeMS04 (Flow) HZMetro KnowAir
-RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE
-STGCN 35.55 22.70 14.59 34.85 21.33 13.47 20.64 15.27 12.92
-STSGCN 33.65 21.19 13.90 33.24 20.79 13.06 20.25 15.22 12.69
-STG-NCDE 31.09 19.21 12.76 32.91 20.75 12.88 20.08 14.93 11.87
-DDGCRN 30.51 18.45 12.19 31.69 19.52 12.45 19.63 12.45 11.84
-ASTGCN 35.22 22.93 16.56 34.36 21.12 13.50 20.37 15.07 12.11
-ASTGNN 31.16 19.26 12.65 32.75 20.63 12.47 19.96 12.38 11.23
-PDFormer 29.97 18.32 12.10 30.18 19.13 11.92 18.65 12.21 11.06
-STAEformer 30.18 18.22 11.98 29.94 18.85 12.03 18.49 11.93 10.75
-MultiSPANS 30.46 19.07 13.29 30.31 18.97 11.85 18.57 12.24 10.88
-STG-Mamba 29.53 18.09 12.12 29.26 18.38 11.67 18.05 11.73 10.66
+RMSE MAE MAPE $std_{\text{MAE}}$ RMSE MAE MAPE $std_{\text{MAE}}$ RMSE MAE MAPE $std_{\text{MAE}}$
+STGCN 35.55 22.70 14.59 1.2068 34.85 21.33 13.47 1.1835 11.46 8.37 11.26 0.4358
+STSGCN 33.65 21.19 13.90 1.1820 33.24 20.79 13.06 1.1249 11.22 8.15 10.89 0.4134
+STG-NCDE 31.09 19.21 12.76 1.0736 32.91 20.75 12.88 1.1097 10.85 7.93 10.47 0.3922
+DDGCRN 30.51 18.45 12.19 1.0365 31.69 19.52 12.45 1.0586 10.43 7.84 10.38 0.3794
+ASTGCN 35.22 22.93 16.56 1.1853 34.36 21.12 13.50 1.1258 10.27 7.57 10.51 0.3745
+ASTGNN 31.16 19.26 12.65 1.0583 32.75 20.63 12.47 1.0784 9.68 7.35 10.23 0.3569
+PDFormer 29.97 18.32 12.10 1.0217 30.18 19.13 11.92 1.0205 9.46 7.12 10.06 0.3375
+STAEformer 30.18 18.22 11.98 0.9678 29.94 18.85 12.03 0.9728 8.69 6.93 9.89 0.3083
+MultiSPANS 30.46 19.07 13.29 0.9851 30.31 18.97 11.85 0.9931 8.57 6.84 10.05 0.3149
+STG-Mamba 29.53 18.09 12.11 0.9218 29.23 18.26 11.59 0.9271 8.05 6.37 9.64 0.2648
 4.5 Robustness Analysis
-Table 3: Robustness analysis on PeMS04. We compare the forecasting performance among STG-Mamba and 4 superior baseline methods during rush/non rush hours, weekend/non weekend settings.
-Models Rush Hours Non Rush Hours Weekend Non Weekend
-RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE
-ASTGNN 33.57 21.05 13.81 29.76 18.22 12.49 30.66 19.08 13.35 31.82 19.53 12.28
-PDFormer 31.69 19.54 12.38 28.95 18.16 11.97 29.87 18.26 12.04 30.39 18.68 12.23
-STAEformer 31.36 19.10 12.22 29.47 17.98 11.93 29.83 18.19 11.96 30.26 18.42 12.27
-MultiSPAN 30.95 19.28 13.31 30.04 18.37 12.54 30.13 18.42 12.73 30.52 19.31 13.34
-STG-Mamba 29.58 18.13 12.15 29.23 18.04 11.91 29.31 18.05 11.92 29.57 18.15 12.16
-STG data exhibit notable periodicity and diversity.
-For urban mobility/traffic data, there are distinct differences in traffic characteristics between the peak travel periods in the morning and evening and the none-peak periods.
-Moreover, the traffic patterns over the 24 hours of weekdays and those of weekends also differ substantially.
-In light of the variations in STG characteristics due to external environmental changes, it is imperative to assess whether deep learning models can accurately and effectively model the spatial-temporal dependencies under various external conditions.
-As such, we establish four distinct external scenarios: (a) Traffic rush hours: Monday to Friday, 8:00-11:00 AM and 16:00-19:00 PM; (b) Non-rush hours: encompassing the remaining time periods from Monday to Friday; (c) Weekend: including all hours on Saturday and Sunday; and (d) Non-weekend: including all time periods from Monday to Friday.
+Table 3: Robustness analysis. We choose rush/non-rush, weekend/non-weekend traffic scenarios.
+Models Rush Hours Non Rush Hours
+RMSE MAE MAPE RMSE MAE MAPE
+ASTGNN 33.57 21.05 13.81 29.76 18.22 12.49
+PDFormer 31.69 19.54 12.38 28.95 18.16 11.97
+STAEformer 31.36 19.10 12.22 29.47 17.98 11.93
+MultiSPAN 30.95 19.28 13.31 30.04 18.37 12.54
+STG-Mamba 29.58 18.13 12.15 29.23 18.04 11.91
+Models Weekend Non Weekend
+RMSE MAE MAPE RMSE MAE MAPE
+ASTGNN 30.66 19.08 13.35 31.82 19.53 12.28
+PDFormer 29.87 18.26 12.04 30.39 18.68 12.23
+STAEformer 29.83 18.19 11.98 30.26 18.42 12.27
+MultiSPAN 30.13 18.42 12.73 30.52 19.31 13.34
+STG-Mamba 29.31 18.05 11.92 29.57 18.15 12.16
+STG data exhibit notable periodicity and diversity, with urban mobility/traffic data showing clear differences between peak travel times in the morning/evening and non-peak periods, as well as between weekdays and weekends.
+Given these variations due to external environmental changes, it is significant to determine whether deep learning models can effectively model the spatiotemporal dependencies under different conditions.
+As such, we establish four distinct external scenarios: (a) rush hours on weekdays from 8:00-11:00 AM and 4:00-7:00 PM; (b) non-rush hours during weekdays; (c) weekends (all hours); and (d) non-weekend (all hours).
 Extensive experiments were conducted on the PeMS04 dataset.
-Table 3 presents the STG forecasting results corresponding to the four traffic scenarios.
-We compare the prediction performance of STG-Mamba with the other four superior baselines, and report the RMSE/MAE/MAPE metrics.
-An ideal deep learning-based STG system should maintain robustness to external disturbances (changes of external environment or events).
-This is specifically reflected in the fact that when facing different scenarios, the model performance should maintain as little difference as possible, rather than being good at some scenarios but being bad at other occasions.
-Among the compared methods, ASTGNN, PDFormer, and STAEformer exhibit relatively large differences in the prediction results between the Rush hour and Non-rush Hour scenarios.
-In the Weekend/Non Weekend scenario, although the performance differences of each model are not as significant as in the Rush hour/Non-Rush hour scenario, we observe that STG-Mamba still shows the smallest performance difference (the maximum robustness) and achieves the best RMSE/MAE/MAPE metrics.
-In summary, STG-Mamba proves the best robustness and outperforms existing baseline methods even in face of complex and diversity traffic scenes.
+Table 3 presents the forecasting results of the four traffic scenarios.
+An ideal deep learning-based STG system should be robust to external disturbances, maintaining consistent performance across different scenarios.
+Compared methods like ASTGNN, PDFormer, and STAEformer show significant performance differences between Rush hour and Non-rush Hour scenarios.
+Even in the Weekend/Non-Weekend scenario, STG-Mamba demonstrates the smallest performance difference, showing maximum robustness and achieving the best RMSE/MAE/MAPE metrics.
+In summary, STG-Mamba exhibits superior robustness and outperforms existing baselines in diverse traffic conditions.
+Statistical-based evaluation in spatiotemporal forecasting is indispensible as they provide a quantifiable measure of how well a model captures and predicts complex data dynamics over time and space.
+Specifically, $R^{2}$ and $\Delta_{\text{VAR}}$ help in assessing the accuracy and robustness of models under varying conditions, ensuring the reliability of predictions and effectively inform decision-making processes.
+The Statistical evaluation results are presented in Table 4.
+Here, STG-Mamba demonstrates superior performance across datasets, achieving the highest $R^{2}$ values and the lowest $\Delta_{\text{VAR}}$ scores, indicating the advancement of accuracy and efficiency in handling spatiotemporal dependencies.
+Table 4: Statistical evaluation results.
+Model PeMS04 (Flow) HZMetro KnowAir
+$\Delta_{\text{VAR}}$ $R^{2}$ $\Delta_{\text{VAR}}$ $R^{2}$ $\Delta_{\text{VAR}}$ $R^{2}$
+STGCN 0.23145 0.77069 0.25769 0.74385 0.49851 0.50487
+STSGCN 0.20538 0.79456 0.21474 0.78862 0.47536 0.52831
+STG-NCDE 0.18275 0.81973 0.19827 0.80492 0.46173 0.54125
+DDGCRN 0.13291 0.87355 0.14796 0.85511 0.43185 0.56506
+ASTGCN 0.17289 0.82698 0.18131 0.82134 0.42729 0.58461
+ASTGNN 0.14863 0.84850 0.15765 0.84627 0.39726 0.60563
+PDFormer 0.11849 0.88190 0.13259 0.87095 0.36182 0.64719
+STAEformer 0.09631 0.90632 0.10383 0.89738 0.33215 0.66085
+MultiSPANS 0.06218 0.93804 0.07596 0.92561 0.32569 0.67122
+STG-Mamba 0.04387 0.95489 0.05148 0.94752 0.30182 0.69537
 4.6 Ablation Study
 Figure 3: Model component analysis of STG-Mamba.
-To investigate the effectiveness of each model component within STG-Mamba, we further design three kinds of model variants, and evaluate their forecasting performance on PeMS04, HZMetro, and KnowAir dataset.
-The detailed description of these model variants are introduced in the following:
-  1. 1.
-    STG-Mamba: The full STG-Mamba model without any modification.
-  2. 2.
-    STG-Mamba w/o KF-Upgrading: We replace the KF-Upgrading module in KFGN with the simple sum average operation to fuse the three temporal branches.
-  3. 3.
-    STG-Mamba w/o Dynamic Filter Linear: We replace the proposed Dynamic Filter Linear in KFGN with the basic static graph convolution.
-  4. 4.
-    STG-Mamba w/o SSSM: The State Space Selection Mechansim (SSSM) is removed from the ST-S3M module.
-  5. 5.
-    STG-Mamba w/o ST-S3M: The whole ST-S3M module is removed from GS3B Encoder.
+To investigate the effectiveness of each model component within STG-Mamba, we further design five kinds of model variants, and evaluate their forecasting performance on PeMS04/HZMetro/KnowAir dataset: (I) STG-Mamba: The full STG-Mamba model without any modification.
+(II) STG-Mamba w/o KF-Upgrading: We replace the KF-Upgrading module in KFGN with the simple sum average operation to fuse the three temporal branches.
+(III) STG-Mamba w/o Dynamic Filter Linear: We replace the proposed Dynamic Filter Linear in KFGN with the basic static graph convolution.
+(IV) STG-Mamba w/o GSSSM: The Graph State Space Selection Mechansim (GSSSM) is replaced by the basic SSSM Gu and Dao (2023).
+(V) STG-Mamba w/o ST-S3M: The whole ST-S3M module is removed from GS3B Encoder.
 As illustrated in Figure 3, STG-Mamba w/o KF-Upgrading shows worse performance than the full model, demonstrating the necessity and suitability of employing statistical learning-based GNN state space upgrading and optimization for SSSM-based methods.
@@
 We propose STG-Mamba that leverages modern SSSM for accurate and efficient STG forecasting.
-STG-Mamba is formulated following the stacked Residual Encoder architecture, with the ST-S3M module facilitating input-dependent graph evolution and node feature selection, successfully integrates STG network with SSSMs.
+In STG-Mamba, the ST-S3M module facilitates input-dependent graph evolution and feature selection, successfully integrates STG network with SSSMs.
 Through the Kalman Filtering Graph Neural Networks (KFGN), the learned STG embeddings achieve a smooth optimization & upgrading based on statistical learning, aligning with the whole STG selective state space modeling process.
-Compared with Attention-based methods, STG-Mamba is able to achieve linear-time complexity and substantial reduction in FLOPs and test inference time.
-We conduct extensive empirical studies on three open-sourced STG datasets from urban traffic, metro In/Out flow, and weather records.
-Experimental results demonstrate the consistent superiority of STG-Mamba over other benchmark methods.
-We believe that STG-Mamba presents a brand new promising approach to general STG learning fields, offering competitive model performance and input-dependent context-aware feature selection under an affordable computational cost.
+Compared with Attention-based methods, STG-Mamba achieves linear-time complexity and substantial reduction in FLOPs and inference time.
+Extensive empirical studies are conducted on three open-sourced STG datasets, demonstrating the consistent superiority of STG-Mamba over other benchmark methods.
+We believe that STG-Mamba presents a brand new promising approach to general STG learning fields, offering competitive model performance and input-dependent contextual feature selection under an affordable computational cost.
+Appendix A ST-S3M Graph Selective Scan Algorithm & Parameter Calculation and Upgrade Algorithm
+Algorithm 1 Graph Selective Scan Algorithm
+Input: $u,\Delta^{*},\mathbb{A},\mathbb{B},\mathbb{C},\mathbb{D}$;
+Output: $z$;
+1:  Obtain the dim of $u$ as $(b,l,d_{\text{in}})$ the second dim of $\mathbb{A}$ as $n$;
+[…]
+[…]
+[…]
+[…]
+6:  Discretize continuous parameters $\mathbb{A}$ and $\mathbb{B}$:
+7:  $\text{delta}\mathbb{A}\leftarrow\exp(\text{einsum}(\Delta^{{}^{\prime}},\mathbb{A}))$, where $\text{delta}\mathbb{A}\in\mathbb{R}^{b\times l\times d_{\text{in}}\times n}$;
+8:  $\text{delta}\mathbb{B}_{u}\leftarrow\text{einsum}(\Delta^{{}^{\prime}},\mathbb{B},u)$, where $\text{delta}\mathbb{B}_{u}\in\mathbb{R}^{b\times l\times d_{\text{in}}\times n}$;
+9:  Initialize State $x$ as zeroes with dimension $\mathbb{R}^{b\times d_{\text{in}}\times n}$;
+10:  for $i$ from $1$ to $l$ do
+11:     $x\leftarrow\text{delta}\mathbb{A}[:,i]\times x+\text{delta}\mathbb{B}_{u}[:,i]$;
+12:     $z\leftarrow\text{einsum}(x,\mathbb{C}[:,i,:])$;
+13:     Append $z$ to the output list $z_{s}$;
+14:  end for
+15:  $z\leftarrow\text{stack}(z_{s})$, where $z\in\mathbb{R}^{b\times l\times d_{\text{in}}}$;
+16:  Adding the direct gain $\mathbb{D}$:
+17:  $z\leftarrow z+u\times\mathbb{D}$
+18:  Return $z$.
+Algorithm 2 State Space Selection Mechanism Parameter Calculation & Upgrade Algorithm.
+Input: $x\in\mathbb{R}^{b\times l\times d_{\text{model}}}$, learnable model params $\mathbb{A}_{log}\in\mathbb{R}^{d_{\text{in}}\times n}$ and $\mathbb{D}\in\mathbb{R}^{d_{\text{in}}}$;
+Output: $y\in\mathbb{R}^{b\times l\times d_{\text{inner}}}$;
+1:  Compute input-independent params $\mathbb{A}\in\mathbb{R}^{d_{\text{in}}\times n}$ and $\mathbb{D}\in\mathbb{R}^{d_{\text{in}}}$:
+2:  $\mathbb{A}=-\exp(\mathbb{A}_{log})$, $\mathbb{D}=\mathbb{D}$;
+3:  Process input $x$ with Linear Layer $Input\_proj$:
+4:  $x_{\text{dbl}}=Input\_proj(x)$, $x_{\text{dbl}}\in\mathbb{R}^{b\times l\times(dt_{\text{rank}}+2n)}$;
+5:  Dividing $x_{\text{dbl}}$ in order to get the input-dependent params $\Delta,\mathbb{B},\mathbb{C}$, where $\Delta\in\mathbb{R}^{b\times l\times dt_{\text{rank}}}$, $\mathbb{B},\mathbb{C}\in\mathbb{R}^{b\times l\times n}$;
+6:  Use Linear Layer $dt\_proj$ and $SoftPlus$ activation to adjust $\Delta$:
+7:  $\Delta^{*}=SoftPlus(dt\_proj(\Delta))$, where $\Delta^{*}\in\mathbb{R}^{b\times l\times d_{\text{in}}}$;
+8:  Employ the GraphSelectiveScan algorithm to proceed with $x,\Delta^{*},\mathbb{A},\mathbb{B},\mathbb{C},\mathbb{D}$:
+9:  $y=\text{GraphSelectiveScan}(x,\Delta^{*},\mathbb{A},\mathbb{B},\mathbb{C},\mathbb{D})$;
+10:  Return the Output $y$.
+Appendix B Computational Efficiency Analysis and Comparison
+B.1 Theoretical Computational Efficiency Analysis
+Leveraging a data-specific selection process, the increasing of sequence length $L$ or graph network node scale $N$ by $k$-fold can results in huge expansion of model parameters for Transformer-based methods.
+STG-Mamba introduces a sophisticated algorithm that is conscious of hardware capabilities, exploiting the layered structure of GPU memory to minimize the ensuing cost.
+In detail, for a batch size of $B$, STG feature dimension of $d^{\prime}$, and network node number of $N$, STG-Mamba reads the $O(BLd^{\prime}+Nd^{\prime})$ input data of $\mathbb{A},\mathbb{B},\mathbb{C},\Delta$ from High Bandwidth Memory (HBM), processes the intermediate stages sized $O(BLd^{\prime}N)$ in Static Random-Access Memory (SRAM), and dispatches the ultimate output, sized $O(BLd^{\prime})$, back to HBM.
+In contrast, if a Transformer-based model is employed, the HBM will read $O(BL^{2}d^{\prime})$ input data, and the computational size remains $O(BL^{2}d^{\prime})$ in SRAM.
+In common practices we often have $BL^{2}d\gg BLd^{\prime}+Nd^{\prime}$, which means SSSM-based approach can effectively decrease the computational overhead in large-scale parallel processing, such as neural network training.
+Additionally, by not retaining intermediate states, it also diminishes memory usage, since these states are recalculated during the gradient assessment in reverse phase.
+SSSM’s GPU-optimized design ensures STG-Mamba operates with a linear time complexity $O(L)$ relative to the length of the input sequence, marking a substantial speed advantage over traditional transformers’ dense attention computations, which exhibit quadratic time complexity $O(L^{2})$.
+B.2 Experimental Computational Efficiency Analysis
+STG-Mamba offers substantial improvement in computational efficiency in addition to performance gain.
+In order to quantatively evaluate computational efficiency, we select the widely recognized Floating Point Operations (FLOPs) and Inference Time as criteria.
+For both of the two criteria, a smaller value indicates better model performance.
+STG-Mamba is compare with other Transformer-based benchmark methods: ASTGNN, PDFormer, and STAEformer, on the same testing set.
+Table 5 showcases the inference time comparison on PeMS04 and KnowAir testing set, we report the average inference time of every 100 batch with Sequence_Length=12, Batch_Size=48.
+It can be seen that ASTGNN is relatively redundancy, with the lowest inference time among the compared Transformer-based methods.
+STG-Mamba gains 18.09% and 20.54% performance increase in inference time than the 2nd best performance model (STAEformer).
+We further compare the computational FLOPs increment with accordance to the scale complexity of STG network.
+Specifically, we use the number of node individuals to represent the scale of STG network.
+In the experiments, we select different number of nodes scales: $[50,100,150,200,250,300]$ for PeMS04 and $[30,60,90,120,150,180]$ for KnowAir, respectively.
+STG-Mamba is compared with the benchmark STAEformer in every settings.
+Figure 4 visualizes the FLOPs evaluation under different STG scale complexity.
+We can easily find that STG-Mamba enjoys a linear increment in FLOPs, while STAEformer shows quadratic increment in FLOPs.
+At the initial small-scale setting, the FLOPs for STG-Mamba/STAEformer are quite similar.
+However, with the increasing of network complexity and scale, the difference becomes more significant.
+At the 300 node setting of PeMS04, the FLOPs for STG-Mamba and STAEformer is 60.85G and 134.92G (2.22 times more), respectively.
+At the 180 node setting of KnowAir, the FLOPs for STG-Mamba and STAEformer is 34.03G and 102.79G (3.54 times more), respectively.
+Benefiting from the linear computational increment brought by modern selective state space models, STG-Mamba and alike models demonstrate stronger capability and computational efficiency for modeling large-scale STG data, which holds promise for addressing the computational overhead in future large-scale STG learning.
+Table 5: Test inference time comparison with benchmark Transformer-based methods.
+Model Inference Time on PeMS04 Inference Time on KnowAir
+ASTGNN 3.285 s 3.018 s
+PDFormer 2.431 s 2.357 s
+STAEformer 2.106 s 1.933 s
+STG-Mamba 1.547 s 1.481 s
+performance $\uparrow$ +26.54% +23.38 %
+Figure 4: (Computational Efficiency Eval) FLOPs comparison between STG-Mamba and STAEformer on PeMS04/KnowAir dataset under different STG node number settings.
+Appendix C Evaluation Metrics Calculation Method
+We employ RMSE/MAE/MAPE/$std_{\text{MAE}}$ as the deterministic evaluation metrics, all the indicators are smaller the better; $\rm{\Delta_{\text{VAR}}}$ (Change in Variance Accounted For) and $\rm{R^{2}}$ (Coefficient of Determination) as statistic-based evaluation metric, where $\rm{\Delta_{\text{VAR}}}$ is the smaller the better, $\rm{R^{2}}$ is the larger the better.
+The reported results in Experiment section are the average value of 10 runs.
+Here, $std_{\text{MAE}}$ refers to the standard deviation of multiple MAE values in the testing stage.
+$std_{\text{MAE}}$ provides a measure of consistency for MAE results across multiple evaluation sets.
+A smaller $std_{\text{MAE}}$ indicates the model performs more stably under different conditions, with less variation in prediction errors.
+The calculation formula for $std_{\text{MAE}}$ is as follows: Step1-Calculate the average MAE $\mu$: $\mu=\frac{1}{n}\sum_{i=1}^{n}\text{MAE}_{i}$.
+Here, $n$ denotes the number of batches in the testing dataset, with the MAE value computed for each batch.
+$\text{MAE}_{i}$ is the MAE result of the $i$-th batch.
+Step2-Calculate the standard deviation: $\text{std}_{\text{MAE}}=\sqrt{\frac{\sum_{i=1}^{n}(\text{MAE}_{i}-\mu)^{2}}{n}}$.
+$R^{2}$ refers to the Coefficient of Determination is a statistical index used for measuring the proportion of variance in the dependent variable that is predictable from the independent variables.
+$R^{2}$ provides a gauge of the effectiveness of the model in explaining the variations in the observed data.
+A higher $R^{2}$ value indicates the model has a higher predictive accuracy and explains a substantial proportion of the variance in the dependent variable.
+$R^{2}$ is formulated as follows: $R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ where $y_{i}$ denotes the ground-truth values, $\hat{y}_{i}$ denotes the model predictions, $\bar{y}$ is the mean of the ground-truth values.
+$\sum(y_{i}-\hat{y}_{i})^{2}$ indicates the unexplained variance by the model, while the denominator, $\sum(y_{i}-\bar{y})^{2}$, is the total variance in the observed data.
+Change in Variance Accounted For ($\Delta_{VAR}$) quantifies the improvement or regression in how much variance the model can account for compared to a previous models or baselines.
+This metric is particularly valuable in iterative model optimization where understanding the variance captured by model updates is crucial.
+$\Delta_{VAR}$ is calculates as follows: $\Delta_{VAR}=\frac{\text{Var}(y-\hat{y})}{\text{Var}(y)}$.
+$\text{Var}(y-\hat{y})$ computes the variance of residuals, and $\text{Var}(y)$ is the variance of the ground-truth data.
+The difference between the variance ratio of the current and previous models provides the change in variance accounted for, reflecting the model’s improvement in explaining the data variability.
+Appendix D Formulation of Selective State Space Models for STG Learning
+SSSMs like structured state space sequential models (S4) and Mamba, initially designed for sequential data, are adapted for Spatial-Temporal Graph (STG) data modeling by leveraging their structure to handle spatiotemporal dependencies dynamically.
+The adaptation involves discretizing the continuous system described by linear Ordinary Differential Equations (ODEs) into a format amenable to deep neural networks modeling and optimization, enhancing computational efficiency while capturing the evolving dynamics of the STG networks.
+Modern selective state space models, such as Mamba, build upon classical continuous systems by incorporating deep learning techniques and discretization methods suitable for complex time-series data.
+Mamba specifically maps a one-dimensional input sequence $x(t)\in\mathbb{R}$ through intermediate implicit states $h(t)\in\mathbb{R}^{N}$, culminating in an output $y(t)\in\mathbb{R}$.
+The fundamental process is described by a modified set of state space equations that include discretization adjustments to accommodate deep learning frameworks:
+$\begin{array}[]{lr}h^{\prime}(t)=\overline{\mathbb{A}}h(t)+\overline{\mathbb{B}}x(t)\\ y(t)=\mathbb{C}h(t)\end{array}$ (17)
+where $\overline{\mathbb{A}}\in\mathbb{R}^{N\times N}$ and $\overline{\mathbb{B}}\in\mathbb{R}^{N\times 1}$ are discretized representations of the continuous state transition and input matrices, respectively.
+$\mathbb{C}\in\mathbb{R}^{1\times N}$ is the output matrices.
+The discretization process involves transformation techniques that convert $\mathbb{A}$ and $\mathbb{B}$ from their continuous forms to computational efficient formats.
+These transformations are determined by a selective scan algorithm that utilizes a convolutional approach to model interactions across temporal dimension:
+$\begin{array}[]{lr}\overline{\mathbb{A}}=\exp(\Delta\mathbb{A})\\ \overline{\mathbb{B}}=\Delta\mathbb{B}\cdot(\exp(\Delta\mathbb{A})-\mathbb{I})\end{array}$ (18)
+Mamba extends classic state space model functionalities by incorporating a selective mechanism that dynamically adapts to the input sequence characteristics.
+This adaptability is achieved through a novel architecture that selectively applies convolution kernels across the sequence, optimizing the model’s response to spatiotemporal variations:
+$\begin{array}[]{lr}\overline{\mathbb{K}}=(\mathbb{C}\overline{\mathbb{B}},\mathbb{C}\overline{\mathbb{A}\mathbb{B}},\dots,\mathbb{C}\overline{\mathbb{A}^{L-1}\mathbb{B}})\end{array}$ (19)
+where $\overline{\mathbb{K}}$ denotes a structured convolutional kernel, capturing dependencies across spatial and temporal dimensions, and $L$ represents the length of the input sequence $x$.
+This selective convolutional approach significantly enhances the model’s capability to forecast spatiotemporal data by focusing computation on critical features of the data sequence, thus improving both accuracy and efficiency in predictions.
+Appendix E Discussion of Limitations
+STG-Mamba for spatial-temporal graph learning demonstrates significant advantages, such as superior forecasting accuracy, model robustness, computational efficiency, and the ability to effectively handle datasets with noise and uncertainty.
+Despite these promising results, there are some limitations to consider:
+  • Although STG-Mamba has been extensively evaluated on three diverse open-sourced datasets (Weather/Traffic/Urban Metro), these datasets may not fully representing the wide range of possible spatiotemporal scenarios, potentially limiting the generalizability of the model.
+Future work should aim to test the model on a broader range of datasets from different geographical regions and temporal scales.
+  • The model interpretability is another concern worthy of further investigation.
+Same as other deep learning-based methods, the black-box nature of Mamba architecture makes it not easy to interpret how specific predictions are made.
+Enhancing the interpretability deep learning-based SSSMs could be beneficial, particularly in applications where understanding the decision-making process is crucial.
+  • While the novel integration of Kalman Filtering statistic learning-based optimization with dynamic graph neural networks helps in managing noise and uncertainty, the model’s performance can still be affected by the quality and completeness of input dataset.
+Developing robust data pre-processing techniques and improving the model’s ability to directly handle incomplete or noisy data would further enhance its reliability.
+We believe by addressing the aforementioned limitations, the future iterations of STG-Mamba can achieve even greater accuracy, robustness, and applicability across a broader range of spatial-temporal graph learning tasks.
+Appendix F Illustration of the STG feature selection procedure
+Figure 5 is the illustration of the spatial-temporal state space selection mechanism that facilitates input-dependent feature selection and focusing.
+Figure 5: Illustration of the spatial-temporal state space selection mechanism that facilitates input-dependent feature selection and focusing.
+Appendix G Related Works
+Owning to the strong ability of capturing dynamic features across both temporal and spatial dimension, attention-based methods, such as Spatial-Temporal Transformer and its variants have recently been broadly implemented in spatial-temporal graph (STG) learning tasks.
+In the domain of STG traffic forecasting, seminal works have emerged that leverage these methods to enhance prediction accuracy.
+ASTGCN Guo et al.
+[2019] introduces an innovative approach combining spatial-attention & temporal attention mechanisms with graph neural networks.
+ADCT-Net Kong et al.
+[2024] presents to incorporate a dual-graphic cross-fusion Transformer for dynamic traffic forecasting.
+Huo et al.
+[2023] integrates hierarchical graph convolutional networks with Transformer networks to capture sophisticated STG relationships among traffic data.
+For weather forecasting, Liang et al.
+[2023] exemplifies how Transformer-based models can be utilized for accurate air quality forecasting across vast geographical areas.
+Chen et al.
+[2023] further demonstrates the power of GNNs in predicting air quality, emphasizing the importance of group-aware mechanisms in enhancing STG forecasting performance.
+In the realm of Social Recommendation, Zhang et al.
+[2023] introduces an Adaptive Disentangled Transformer (ADT) framework that optimizes the disentanglement of attention heads in different layers of Transformer architecture.
+Wei et al.
+[2023] explores the improved efficiency of multimedia recommendation by introducing modal-specific embeddings and a lightweight self-attention.
+Although Transformer-based methods demonstrate notable improvements in STG relationship learning Chen et al.
+[2022], Cong et al.
+[2021], their modeling ability on large-scale STG networks and Long-Term Sequential Forecasting (LTSF) tasks is greatly hindered by the quadratic computational complexity $O(n^{2})$ provided by its core component–attention mechanism Zhou et al.
+[2021, 2022], Rampášek et al.
+[2022].
+Moreover, attention mechanisms usually approximate the full-length historical data, or encode all the context features, which may not be necessary for modeling the long-term dependencies of STG learning tasks.
+As such, the research world has been actively looking for counterparts of Attention-based models.